### v2

Abstract
Mixture-of-Experts (MoE) has emerged as a practical approach to scale up parameters for the Transformer model to achieve better generalization while maintaining a sub-linear increase in computation overhead. Current MoE models are mainly built with expert parallelism on distributed devices. However, it usually depends on homogeneous devices to deploy and suffers from heavy communication overhead and computation redundancy. In this paper, we explore developing a Heterogeneous-aware EXpert Allocation framework, HEXA-MoE, with significantly enhanced computing efficiency. It contains two components: ($1$) Expert-Specific Operators. We replace the typical general matrix multiplication or grouped matrix multiplication interfaces with our operators, which allows the computing to be performed in an in-place manner with ZERO redundancy. ($2$) Adaptive Data- and Model-Centric Configurations for different workload scales. Specifically, we introduce a pipeline-shared cache on each device to tackle the heavy memory consumption in the existing data-centric MoE library. Comprehensive experiments on the Swin-MoE benchmark consistently reveal the effectiveness of our HEXA-MoE framework, i.e., reducing $10\%\sim 48\%$ memory consumption and achieving 0.54.3×0.5$\sim$ 4.3$\times$0.5 ∼ 4.3 × speed up compared to current state-of-the-art MoE libraries. Furthermore, we examine our HEXA-MoE with heterogeneous devices for both data- and model-centric settings. Promising results show that employing optimal parallel configuration with HEXA-MoE on heterogeneous devices can substantially minimize overall latency. Codes are available at here.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Transformers have become the de facto architecture for a vast range of machine learning tasks including natural language process Vaswani (2017); Raffel et al. (2020), computer vision Liu et al. (2021); He et al. (2022b) and multi-modal learning Li et al. (2022); Liu et al. (2024). To scale up model parameters for stronger learning capacity and better generalization following the scaling law Kaplan et al. (2020), Mixture-of-Experts (MoE) Jiang et al. (2024) has been demonstrated as a practical and computation-efficient approach. For a single MoE layer, the feed-forward network (FFN) is expanded to a group of experts compared to the dense counterpart, activated selectively during computing. This dynamic nature poses peculiar challenges to system design. Current MoE frameworks usually present some unique attributes that differentiate them from dense models:
  1. Distributed computing is required for MoE due to its sheer size, and expert parallelism Lepikhin et al. (2020) is the most commonly used technique, which distributes the experts in one layer on different devices.
  2. Since the workload varies dynamically for each expert in each layer, we have to re-organize the local data batches for expert computing via dispatch and combine operations. Specifically, we have to calibrate the workload for each expert to utilize the general matrix multiplication (GeMM) or grouped GeMM interface.
  3. Since each token may activate any one or several experts and different experts may be distributed on different devices in expert parallelism, it is necessary to communicate among devices before and after each MoE layer via all-to-all communication.
Figure 2: Tensor Parallelism vs Expert Parallelism for MoE layer. We take $4$ experts on $2$ devices as an example and visualize the parameter distribution of $2$ MLPs. For expert parallelism, each device can only keep several experts, while for tensor parallelism the parameter allocation is more flexible since the partition of parameters is along the hidden dimension.
These peculiar designs lead to inefficiency for both training and inference in MoE computing. On the one hand, dispatch and combine operations result in unnecessary token padding or discarding Gale et al. (2023), leading to both redundant computation and compromise on model performance. Meanwhile, conventional expert parallelism depends on all-to-all communication, which requires synchronization and occupies a lot of latency Li et al. (2023). On the other hand, current expert parallelism for MoE models is mainly deployed on homogeneous devices. However, a cluster equipped with the latest homogeneous devices is usually much more expensive than heterogeneous devices due to the iteration of GPU development and the decrease on prices for old-version GPUs. If we want to adapt expert parallelism to heterogeneous devices, it would be necessary to rearrange the expert distribution to utilize the total computing capacity better. However, the inherent dynamic property of the MoE layer makes it difficult to implement, as the workload of each expert changes dynamically in each step during computing He et al. (2022a); Li et al. (2023). Meanwhile, since heterogeneous devices are cheaper and easier to access if MoE computing can be re-designed to be heterogeneous-aware, it would benefit wider deployment.
This paper proposes HEXA-MoE, a heterogeneous-aware framework for efficient MoE computing. We first find that if compiling the MoE layer into GeMM like other operators such as Linear, Convolution, etc., the unnecessary components such as token padding or discarding are unavoidable. To tackle it, we investigate the forward and backward propagation for the MoE layer thoroughly, and propose to replace current GeMM or grouped GeMM Gale et al. (2023) interfaces with expert-specific operators. Specifically, all the computations related to the MoE layer can be summarized as built on 3 operators, i.e., expert-specific matrix multiplication (ESMM), expert-specific summation (ESS) and expert-specific transposed matrix multiplication (ESTMM). The forward propagation only requires ESMM, while backward requires all the $3$ operators. We implement them in a highly-parallelized manner on GPU, and distribute the MoE layer with tensor parallelism Narayanan et al. (2021) as an alternative to expert parallelism, shown in Figure 2.
HEXA-MoE also considers different workload scales for MoE computing and is adapted to both data- and model-centric configurations. Specifically, the difference lies in the content of communication. For model-centric, local data batches are synchronized among devices, and model parameters are kept still, while for data-centric, local parameters are gathered to each device, while local data batches are kept still Liu et al. (2023). When the scale of data is larger than the model parameters, the data-centric setting would perform better. However, previous work preserves all the gathered MoE parameters on each device for backward propagation, leading to huge memory consumption. To achieve better memory efficiency while maintaining high computing speed, we propose to introduce a pipeline-shared cache on each device, which is updated dynamically during both forward and backward propagation, and the all-gather communication can be overlapped with non-MoE operations such as attention. Essentially, a data-centric setting can be viewed as data parallelism, which is easy to adapt to heterogeneous devices. A model-centric setting is employed when the size of model parameters is larger than the data workload. In this case, all gathered communication is applied to the data batches, and tensor parallelism is applied to each MoE layer. Each device processes all the data batches using the parameter chunk it preserves. This can also be employed in heterogeneous devices via adaptively allocating expert parameters, as illustrated in Figure 2.
Our contributions can be summarized as follows:
  1. $\star$
    We propose to refactor MoE computing with expert-specific operators, enabling it to be performed in an in-place manner without the requirement of dispatch and combine operations. In this way, ZERO computation redundancy would be introduced. We implement these operators in a highly parallelized manner on GPU.
  2. $\star$
    We develop HEXA-MoE based on the expert-specific operators for both data- and model-centric settings, considering different scales of workload, and propose to employ pipeline-shared cache for data-centric setting to improve memory efficiency. Experiments with Swin-MoE show that HEXA-MoE can reduce $10\%\sim 48\%$ memory usage and achieve 0.54.3×0.5$\sim$ 4.3$\times$0.5 ∼ 4.3 × speed up compared to state-of-the-art MoE computing libraries.
  3. $\star$
    Our HEXA-MoE transforms conventional expert parallelism into pure data or tensor parallelism, making it heterogeneous-aware. Experiments show that our HEXA-MoE can be effectively adapted to heterogeneous devices, and employing the optimal parallel configuration can substantially minimize overall latency.
Table 1: Description of symbols.
Symbol Description
$E$ Number of global experts.
$N$ Number of the input tokens.
$\mathcal{F}$ Activation function between the two MLPs.
$\mathcal{F}^{\prime}$ Element-wise differential for $\mathcal{F}$.
$\odot$ Element-wise product.
$D_{i},D_{o}$ Input and output size of the FFN.
$\bm{x}$ Data batch with $N$ tokens.
$\bm{x}^{e},N_{e}$ Tokens in $\bm{x}$ routed to expert $e$ with number $N_{e}$.
$\{\mathcal{R}_{i}(\bm{x})\}_{i=0}^{k-1}$ Routing choice for $\bm{x}$ under top-$k$ routing.
$\bm{W_{1}},\bm{W_{2}}$ Weights for 2 MLPs, shaped as $(E,D_{1},D_{2})$.
$\bm{b_{1}},\bm{b_{2}}$ Biases of the two MLPs, shaped as $(E,D)$.
BLK Block size for expert-specific operators.
WARP Warp size in a thread block.
TIMES A thread block has WARP×$\times$×TIMES threads.
2 Related Works
Scaling Transformers with MoE.
Transformer models can be scaled up via the Mixture-of-Experts (MoE) for better learning capacity and generalization while enjoying a sub-linear increase in computation overhead due to its sparsity. It has been proven in many research fields such as natural language processing Fedus et al. (2022); Jiang et al. (2024); Du et al. (2022), machine vision Riquelme et al. (2021); Fan et al. (2022) and multi-modal learning Bao et al. (2022).
Open-Source MoE Computing Libraries.
FastMoE He et al. (2021) pioneered the first PyTorch open-source implementation for distributed MoE. Based on it, FasterMoE He et al. (2022a) proposes dynamic shadowing and smart scheduling to tackle load imbalance and improve parallelism. Tutel Hwang et al. (2023) implements switchable parallelism and dynamic pipelining, improving adaptability and scalability. MegaBlocks Gale et al. (2023) proposes block-sparse operations and corresponding GPU kernels to mitigate dynamic routing issues in MoEs.
Improving efficiency for MoE computing.
MoE computing faces significant challenges in communication and memory. Recent research has explored alleviating them from various perspectives: Lina Li et al. (2023) proposes to dynamically schedule resources to reduce latency and tackle the all-to-all bottleneck. Janus Liu et al. (2023) proposes a data-centric paradigm that replaces traditional all-to-all communication with asynchronous expert fetching, enabling overlapped communication and computation for enhanced parallelism. ScheMoE Shi et al. (2024) proposes a generic scheduling framework for optimal communication and computation scheduling during MoE training. MPMoE Zhang et al. (2024) accelerates MoE training through adaptive and memory-efficient pipeline parallelism. SmartMoE Zhai et al. (2023) proposes an efficient searching algorithm to identify optimization opportunities within an expanded hybrid parallelism space, tailored for data-sensitive MoE models. PipeMoE Shi et al. (2023) adaptively pipelines communications and computations in MoE to mask communication latency. It also provides an optimal strategy to determine pipeline degree to minimize overall iteration time.
3 Preliminary
Figure 4: Thread hierarchy for CUDA programming and memory hierarchy for a typical NVIDIA GPU.
Introduction to CUDA Programming:
A GPU program can be viewed as a thread grid in Figure LABEL:fig:thread-hierarchy, which is $3$-dimensional and contains massive thread blocks. A thread block is also $3$-dimensional and contains up to $1024$ threads. Computation is mapped to these threads and implemented in parallel. Physically a GPU is composed of massive streaming multiprocessors (SM), where we show the memory hierarchy in Figure LABEL:fig:memory-hierarchy. Each SM loads data from global memory to its registers for parallel computing, and caching with shared memory can be faster. An SM contains many processing units including CUDA core and Tensor core. The former is used for general parallel computing, while the latter is specialized for matrix multiplication with mixed precision, which is first introduced in NVIDIA Volta architecture Choquette et al. (2018). During execution, an SM is given one or more thread blocks, which are partitioned into warps and each warp gets scheduled by a warp scheduler.
Mathematical Symbols:
We provide the description of mathematical symbols used in this paper in Table 1, including variables in MoE computing and constant values in CUDA programming. To utilize Tensor core for faster matrix processing, we have to call the nvcuda::wmma (warped matrix multiplication and add) interface to compute a fixed-sized matrix multiplication such as $\times 16$ in a single warp, therefore the number of threads in a block has to be divisible by the constant value WARP. In the CUDA implementation of our method, a thread block loads BLK tokens each time from global memory to shared memory to conduct the parallelized computation.
Figure 5: Comparison between conventional and expert-specific formulation for MoE computing. We take top-1 routing for illustration and present the corresponding relation of each formula in the MoE forward and backward propagation.
4 Method
We start our exploration by first formulating the forward and backward propagation of a single MoE layer with and without GeMM, which induces the same calculation results in different ways (Sec. 4.1). Specifically, GeMM-based MoE computing depends on inefficient dispatch & combine operations, while our expert-specific operators serve as an alternative, performing in-place and inducing ZERO computation redundancy. After that, we expound the details for our optimized CUDA implementation (Sec. 4.2). We also introduce a novel expert-specific fused operator in this part, specialized for parallel MoE backward propagation. These empower our HEXA-MoE with high computation efficiency. Next, we consider different workload scales, and adapt HEXA-MoE efficiently to both data- and model-centric settings (Sec. 4.3). Finally, we adapt HEXA-MoE to heterogeneous devices, and provide an expert allocation approach to minimize the average latency so as to utilize better heterogeneous computing capacities (Sec. 4.4).
4.1 MoE Computing Formulation
GeMM-based MoE Computing.
Current MoE libraries are mainly built on expert parallelism Fedus et al. (2022) and dispatch & combine operations Hwang et al. (2023). Specifically, “expert parallelism” means distributing the experts to different devices, “dispatch” means assigning the experts with different tokens based on the routing choice, and “combine” means restoring the original data batch from dispatched tokens. In conventional settings, all-to-all communication is required to further assign the dispatched tokens to the corresponding device, where the corresponding expert is kept. To employ GeMM on each device, we need to calibrate the workload of each expert via token padding or discarding and build a new data batch. Taking top-1 routing as an example, we formulate the forward and backward propagation in Figure 5, denoting $\ell$ as loss value and $\bm{x}^{e}$ as tokens dispatched to expert $e$ (with $N_{\textrm{e}}$ tokens). In backward propagation, the gradients of the output have been provided by the auto-differentiation program.
Based on dispatch and combine operations, all variables can be derived via basic matrix operations such as summation and multiplication. Although we can utilize the high-performance GeMM interface, these operations are inefficient, since the workload of each expert varies dynamically in each step, and token padding or discarding has to be employed to construct new data batches for local experts. It can be overcome by our expert-specific design.
MoE Computing via Expert-Specific Operators.
Based on the above formulations, we propose to re-implement in an in-place manner to address the inefficient teaser of dispatch & combine operations. Specifically, we find that the forward and backward propagation of a single MoE layer can be reformulated as built upon 3 specialized operators from an expert-specific perspective, namely expert-specific operators. We take top-$1$ routing for illustration, while formulations for top-$k$ routing are provided in Appendix A.
Figure 9: Illustration of the proposed operators. Take $10$ tokens, $4$ global experts, and block size $4$ as an example. For ESTMM, the $2$ input batches are in a re-indexed format, while for others both the raw batch and re-index vector are provided.
  1. Expert-specific matrix multiplication (ESMM): given input batch $\bm{x}$ with $N$ tokens, routing choice $\mathcal{R}(\bm{x})$, weight $\bm{W}$ and bias $\bm{b}$, we formulate it as $\bm{y}=\textit{ESMM}(\bm{x},\bm{W},\bm{b},\mathcal{R}(\bm{x}))$, where $\bm{y}_{i}$ is derived from $\bm{x}_{i}$, as well as the corresponding $\bm{W}_{\mathcal{R}(\bm{x}_{i})}$ and $\bm{b}_{\mathcal{R}(\bm{x}_{i})}$. ESMM is applied to get the hidden and output tokens in forward, and the gradients of hidden and input tokens in backward propagation, as shown in Figure LABEL:fig:esmm.
  2. Expert-specific summation (ESS): given input batch $\bm{x}$ with $N$ tokens and routing choice $\mathcal{R}(\bm{x})$, it is formulated as $\bm{y}=\textit{ESS}(\bm{x},\mathcal{R}(\bm{x}))$, where tokens routed to expert $e$ are accumulated and recorded in $\bm{y}[e]$. ESS is applied to get the gradients of bias in backward propagation, as illustrated in Figure LABEL:fig:ess.
  3. Expert-specific transposed matrix multiplication (ESTMM): the inputs include two batches $\bm{x}_{1}$ and $\bm{x}_{2}$ (both with $N$ tokens), and the routing choice $\mathcal{R}(\bm{x})$. Both $\bm{x}_{1}$ and $\bm{x}_{2}$ are the ESMM result of $\bm{x}$, thus sharing the same routing choice. It is formulated as $\bm{y}=\textit{ESTMM}(\bm{x}_{1},\bm{x}_{2},\mathcal{R}(\bm{x}))$. For the $i$-th channel of $\bm{x}_{1}$ and $j$-th channel of $\bm{x}_{2}$, we first prepare a zero vector $\bm{c}$ with $E$ elements, and accumulate $\bm{x}_{1}[m,i]\cdot\bm{x}_{2}[m,j]$ to $\bm{c}[\mathcal{R}(\bm{x}_{m})]$ for all $0\leq m<N$. After that we assign $\bm{c}[k]$ to $\bm{y}[k,i,j]$ for all $0\leq k<E$. ESTMM is applied to get the gradients of weights in backward propagation, as illustrated in Figure LABEL:fig:estmm.
Based on these expert-specific operators, we can now implement MoE computing in a novel in-place manner, as formulated in Figure 5. We compare each stage between our method and conventional formulation, and the output of each token is ensured to be consistent for both cases. In forward propagation, ESMM performs as an alternative to common matrix multiplication, while in backward propagation, the gradient of $\bm{y}$ is also provided by the auto-differentiation program. ESS performs as an alternative to tensor summation, while ESMM and ESTMM are designed as $2$ specialized cases for matrix multiplication from the expert-specific perspective. With the help of our expert-specific design, we can now directly compute the output for all MoE computing tasks rather than dispatching the tokens to each expert before MoE computing, introducing zero redundancy.
4.2 Optimized CUDA Implementations
Figure 13: Naive expert-specific MoE computing and the memory-efficient version. We take top-2 routing and 4 tokens as an example. In the naive version, we have to compute from top-$1$ to top-$k$ respectively, write each result back, and sum them up for the final output, while a memory-efficient version adopts AtomicAdd instead, saving much memory. We also present a comparison for the 2 schemes, assuming the memory for one input & output token as 1 and hidden token as 4.
Re-Index based Expert-Specific Operators.
Although expert-specific operators enable MoE computing to be performed in an in-place manner, we should not implement them directly, since simply computing for each token in parallel cannot fully utilize the locality of GPU, which restricts the performance. Considering that we cannot guarantee the adjacent tokens to be routed to the same expert, we have to introduce other tensors as auxiliarys. In HEXA-MoE, we additionally adopt a re-index vector, as illustrated in Figure LABEL:fig:re-index. Specifically, we re-organize the token indices based on the routing choice, i.e., gathering indices for tokens that routed to the same expert into a sub-vector, and padding it with -1 to make it divisible by a constant value BLK. The algorithm details are provided in Appendix A.
We present the details of optimized ESMM implementation in Figure LABEL:fig:esmm, where a thread block loads BLK input tokens each time based on the re-index vector, as well as the corresponding expert parameters. Since these BLK tokens are routed to the same expert, we only need to load one expert each time. We then accumulate the dot product results along the dimension of the input feature and write the result for each token back to its position in global memory, also based on the re-index vector. The details of optimized ESS is shown in Figure LABEL:fig:ess, where each thread block is employed to compute some channels of one expert, with BLK tokens one time. After accumulating all tokens within one expert, it writes the summation back to global memory. ESTMM is illustrated in Figure LABEL:fig:estmm, where we present the input batches in a re-indexed manner. Specifically, the two input batches are the ESMM results of the same tensor, sharing the same re-index vector. Each thread block loads BLK channels of both batches for tokens that are routed to the same expert. After accumulation, results are written back to the corresponding positions respectively. We also provide algorithm details of this part with PyTorch style pseudo-code in Appendix A.
Memory Optimization for Top-$k$ Routing.
We find that when extending the routing choice from top-$1$ to top-$k$, the memory consumption would increase significantly. To tackle this issue, we present a simplified pipeline in Figure 13, where only input, hidden, and output tokens in one MoE layer are provided. For naive forward propagation in Figure LABEL:fig:forward-normal and backward in Figure LABEL:fig:backward-normal, we have to compute the result of all the $k$ routing choices and sum them up to obtain the final output. Specifically, in forward propagation, we have to sum up the outputs of ESMM for the second MLP layer to get the output tokens, while in backward propagation we have to sum up the outputs of ESMM for the first MLP layer to get the gradients for input tokens. This leads to extra memory allocation for pre-summed tensors, including pre-summed output tokens in forward and pre-summed input tokens in backward. To optimize memory consumption when scaling from top-1 routing to top-$k$, we only allocate memory for the hidden tokens, and sum up the result of each routing choice via atomicAdd in CUDA programming, as illustrated in Figure LABEL:fig:memory-efficienct. This can result in obvious memory reduction, as we presented in Figure LABEL:fig:mem-opt.
Figure 15: Visualization of the training pipeline and shared cache under the data-centric setting. Before all gather communication, each device loads the kept parameter chunk to its shared cache, after that each cache will be filled with the whole parameters of an MoE layer.
Figure 16: Tensor parallelism for HEXA-MoE under model-centric setting. Take $2$ GPUs, $1$ MoE layer, and $4$ experts as an example, assuming the local batch contains $4$ tokens. We refactor classical tensor parallelism with expert-specific operators.
Fused Kernel in Backward Propagation.
For a single multi-layer perception (MLP) $\bm{y}=\bm{x}\cdot\bm{W}+\bm{b}$, the gradients for $\bm{x}$, $\bm{W}$ and $\bm{b}$ can be computed in parallel, as they are unrelated with each other. Therefore, we can integrate the implementation of ESS, ESTMM, and ESMM together into one kernel function named expert-specific fused kernel (ESFK) for better parallelism. However, directly integrating them is hard to implement, since the shape of thread blocks required for each operator and the implementation details for each operator vary a lot with each other. To this end, we set the shape of the thread block to be the same for each operator, and expand one dimension for ESMM and ESS so that the thread grids are all $3$-dimensional. Taking a single FFN layer under top-$1$ routing as an example, where we present the shape of allocated thread block and thread grid for both forward and backward propagation in Table 6 in Appendix B. By transposing the shape or expanding one dimension for the thread grid of an operator, we can align the 1st and 2nd dimensions while summing the 3rd dimension together for an integrated grid. In this way, the backward propagation for one FFN can be re-implemented with only $2$ fused kernel function and $1$ element-wise dot production.
4.3 Parallel Strategy for Different Workload Scale
Data Parallelism for Data-Centric Setting.
Data-centric is a practical and efficient approach for both training and inference when the scale of the data batch is larger than the model parameters. HEXA-MoE distributes MoE parameters to different devices along the hidden size dimension of FFN, based on tensor parallelism. In a data-centric setting, each device gathers the whole MoE parameters from other devices and performs computing for each MoE layer locally, similar to data parallelism. Although MoE computing under data-centric configuration has been explored by Janus Liu et al. (2023), memory efficiency has not yet been fully studied. Janus pre-fetchs the required MoE parameters for each layer progressively in forward propagation and preserves all of them locally for backward propagation so that no communication is required when computing for gradients. However, the sheer size of MoE parameters can lead to a significant increase in memory usage, making it inefficient and impracticable for wider deployment.
To tackle this issue, we propose to allocate an additional region on the GPU memory of each device, which is adopted to dynamically cache the gathered MoE parameters for each pipeline stage, named as pipeline-shared cache. Since each device keeps a subset of the FFN hidden dimension for all experts in each layer, we need to employ all gather communication among devices before the computing of each MoE layer, preparing for the required parameters during both forward and backward propagation, as illustrated in Figure LABEL:fig:pipeline and LABEL:fig:cache. To achieve better parallelism, all gather communication can be overlapped with other operations such as attention and router, as shown in Figure LABEL:fig:pipeline. In this way, each device would not have to preserve all the MoE parameters for backward propagation, while communication overhead can also be well masked, therefore both memory efficiency and computing efficiency can be achieved.
Tensor Parallelism for Model-Centric Setting.
In the case that the scale of data is less than model parameters, model-centric configuration turns out to be more efficient than data-centric. To distribute MoE parameters on multiple devices, we modify the classical tensor parallelism with our proposed expert-specific operators, as shown in Figure 16. All gather communication is employed to synchronize local data batches before and after each MoE layer, and a pipeline-shared cache is not required in this setting. In Figure 2 we present a comparison between expert parallelism and tensor parallelism for the MoE layer, where HEXA-MoE takes small modification based on standard tensor parallelism Narayanan et al. (2021). We distribute each MoE layer among different devices along the hidden size dimension for each expert, and during MoE computing, each device computes the all gathered data batches with only the local MoE parameter chunk using ESMM, and the output tokens are all reduced with sum operation in forward propagation. During backward propagation, all gather and all reduce communications are interchanged, while ESMM are replaced by ESMM, ESS, and ESTMM to derive the gradients for input tokens, bias, and weights, respectively, which can also be replaced by the fused operator ESFK.
Figure 18: Average memory usage for training Swin-Transformer-MoE models. We take 8 global experts and examine all cases from top-1 to top-8 routings. Experiments are conducted on 2 homogeneous GPUs using automatic mixed precision in PyTorch. The batch size is set to 40 for all cases. We record the average GPU memory consumption (GB) on each device.
4.4 Heterogeneous-aware Expert Allocation
Workload Division for a different configuration.
For data-centric configuration, since it can be essentially viewed as data parallelism, the workload mainly depends on batch size. For model-centric configuration, the workload of each device for the MoE layer mainly depends on the allocated sub-dimension of the hidden size with each expert, as we have replaced conventional expert parallelism with pure tensor parallelism with the aid of expert-specific operators.
Practical Expert Allocation on Heterogeneous Devices.
To adapt our HEXA-MoE to heterogeneous devices, we propose a specialized expert allocation approach that better utilizes the overall computing capacity by adjusting the workload on each device. Specifically, we have to first examine the computing capacity of each device by averaging its latency on the same proxy task of heavy computing, such as large-scale matrix multiplication or large-scale MLP computing. Denoting the results as $\{t_{i}\}_{i=0}^{N-1}$ for $N$ devices that distribute the MoE model. For the data-centric setting, we assign different local batch size $\{B_{i}\}_{i=0}^{N-1}$ to different devices based on the observed latency, as shown in Equation 1:
$B_{i}=\frac{1/t_{i}}{\sum_{j=0}^{N-1}1/t_{j}}\cdot B_{\textrm{global}},$ (1)
where we denote $B_{\textrm{global}}$ as the global batch size, i.e., $B_{\textrm{global}}=\sum_{j=0}^{N-1}B_{i}$. For the model-centric setting, we denote the sub-dimension of hidden size for one MoE layer on each device as $\{h_{i}\}_{i=0}^{N-1}$ with total hidden size $H$. The assigned sub-dimension $h_{i}$ on device $i$ can be derived as Equation 2.
$h_{i}=\frac{1/t_{i}}{\sum_{j=0}^{N-1}1/t_{j}}\cdot H.$ (2)
To ensure that both $\{B_{i}\}_{i=0}^{N-1}$ and $\{h_{i}\}_{i=0}^{N-1}$ are integers, we need to further round them up or down while making sure that the summation is exactly $B_{\textrm{global}}$ or $H$.
5 Experiments
5.1 Experimental Setup
We implement all the experiments using 2 machines, namely $M_{\textrm{homo}}$ and $M_{\textrm{hete}}$. Specifically, $M_{\textrm{homo}}$ is composed of 4 homogeneous GPUs, while $M_{\textrm{hete}}$ contains 2 heterogeneous GPUs. Details for the machines and GPUs are provided in Table 2. We take the training process of Swin-Transformer-MoE as the benchmark for all the experiments, including memory analysis, latency analysis, heterogeneous computing analysis, and ablation studies. Apart from heterogeneous computing analysis, all the experiments are conducted on $M_{\textrm{homo}}$. We also adopt both the Small and Base scales for the Swin-MoE model, following Tutel Hwang et al. (2023). Experiments are conducted with PyTorch, taking nccl as the communication backend. All the experiments are implemented in a distributed manner based on the distributed data-parallel interface in PyTorch, and batch size is the batch size on a single device rather than the global workload.
Figure 19: Average latency for training Swin-Transformer-MoE models. Experiments are conducted on 4 homogeneous GPUs with 4 global experts. We set different batch sizes for different models under different routing strategies to maximize the utilization of GPU memory. We record the average latency for one step (s) during training with 2k steps in total.
5.2 Experimental Results
Memory Analysis on Homogeneous Devices.
We analyze the memory consumption for different MoE libraries on homogeneous devices. We adopt 8 global experts and examine all cases from top-$1$ routing to top-$8$. Results are visualized in Figure 18. HEXA-MoE can reduce $10\%$-$48\%$ memory usage compared to Tutel Hwang et al. (2023) and MegaBlocks Gale et al. (2023), and the memory usage for model-centric setting is slightly less than data-centric setting due to the employment of pipeline-shared cache. When scaling top-$1$ routing to top-$k$ routing, the memory reduction becomes more significant, since we only enlarge the memory allocation for the hidden tokens, as illustrated in Figure 13. Meanwhile, the trend of memory consumption increase for our method is also much more gentle than Tutel and MegaBlocks when scaling from top-1 routing to top-$k$, owing to the introduction of memory-latency trade-off. Since HEXA-MoE performs MoE computing in an in-place manner and introduces no token padding, we can still save a certain amount of memory under top-$1$ routing.
Table 2: Details of the machines and GPUs used for both homogeneous and heterogeneous experiments.
Notation Number & Hardware Specs Memory
CPU $M_{\textrm{homo}}$ 2×2$\times$2 × Intel Xeon Platinum 8352V 2.10GHz 1008 GB
$M_{\textrm{hete}}$ 2×2$\times$2 × Intel Xeon Gold 6130 2.10GHz 62.5 GB
GPU $D$ (in $M_{\textrm{homo}}$) 4×4$\times$4 × NVIDIA GeForce RTX 4090 24 GB
$D_{0}$ (in $M_{\textrm{hete}}$) 1×1$\times$1 × NVIDIA TITAN RTX 24 GB
$D_{1}$ (in $M_{\textrm{hete}}$) 1×1$\times$1 × NVIDIA GeForce RTX 2080 Ti 11 GB
Latency Analysis on Homogeneous Devices.
We analyze the average latency for one training step, and present the results in Figure 19. To fully utilize the GPU memory, we set different batch sizes for different routing configurations and different model scales. HEXA-MoE achieves $0.5$-$4.3$$\times$ speed up compared to Tutel Hwang et al. (2023) and MegaBlocks Gale et al. (2023). Since the scale of workload is larger than MoE parameters in our experiments, the average latency under a data-centric setting is less than model-centric in these cases, and the advantage of HEXA-MoE turns more obvious while enlarging batch size. To better visualize the latency comparison between data-centric and model-centric configurations, we present the latency for the Swin-MoE-Base model under different batch sizes in Figure 20. When the workload is relatively small, a model-centric setting is more efficient, while a data-centric setting becomes more efficient with a relatively large workload. Experiments are all conducted on the homogeneous machine with automatic mixed precision in PyTorch.
Figure 20: Latency comparison with data-centric and model-centric configurations. We record average latency per step with 2k steps in total. The data-centric setting presents a more gentle trend when scaling workload.
Heterogeneous Experiments.
We demonstrate the heterogeneous awareness of our HEXA-MoE via experiments on different devices under both data-centric and model-centric configurations, and visualize the average latency with different division strategies in Figure 22. Since the $2$ GPUs in our heterogeneous machine have similar computing capacities, we also adjust the power limit for each device. We first examine the computing capacity for these experimental settings, and provide the results in Table 3. The details of the proxy task are provided in Appendix B.
In Figure LABEL:subfig:data-centric-hete, we adjust the batch size on each device under data-centric configuration. Essentially the optimal proportion is very close to the examined computing capacity proportion, and it has a certain deviation from uniform division when the power constraints of the two devices are not equal, while the optimal division is also uniform division when the power constraints are equal. Specifically, when setting the power constraint for $D_{0}$ to be 100W and $D_{1}$ to be 300W, employing our heterogeneous-aware allocation can make the average latency to be 13.2% lower than naive division, and when setting $D_{0}$ to be 300W and $D_{1}$ to be 100W, our approach can reduce the average latency by 25.3%.
In Figure LABEL:subfig:model-centric-hete, we adjust the allocated sub-dimension proportion of the hidden size for each MoE layer on each device. Adapting the sub-dimension proportion to be close to the computing capacity proportion of heterogeneous devices can also substantially minimize average latency, similar to the data-centric setting. Specifically, when setting the power constraint of $D_{0}$ to be 100W and $D_{1}$ to be 300W, our approach can achieve a 6.3% reduction on average latency compared to uniform division, and when setting $D_{0}$ to be 300W and $D_{1}$ to be 100W, we can reduce it by 11.9%. Although the reduction is not as significant as a data-centric setting, we can still achieve a relative speed-up. These demonstrate that our HEXA-MoE can essentially maximize the utilization of heterogeneous computing resources.
Table 3: Examining computing capacity for heterogeneous devices. $P$, $T$ and $R$ denote power constraint (W), average latency (s), and computing capacity proportion.
Case 1 Case 2 Case 3
$P$ $T$ $R$ $P$ $T$ $R$ $P$ $T$ $R$
$D_{0}$ 100 4.58 0.40 300 3.20 0.50 300 3.28 0.74
$D_{1}$ 300 3.06 0.60 300 3.18 0.50 100 9.42 0.26
Figure 22: Latency analysis on heterogeneous devices with both data-centric and model-centric configuration. Experiments are all conducted with Swin-MoE-Small model using 4 global experts and top-2 routing. The average latency falls minimal when the division proportion is close to the computing capacity proportion in each case.
5.3 Ablation Studies
We examine the effectiveness of each component for our HEXA-MoE via measuring the impact of expert-specific operators, pipeline-shared cache, fused kernel, data- & model-centric and memory-latency trade-off respectively on two indicators, i.e., average latency and memory usage. We take the distributed training of the Swin-MoE-Base model on homogeneous devices as the benchmark for ablation study experiments. Since dispatch & combine operations are inevitable if we dispense with our expert-specific operators, we directly take the performance of Tutel in this case.
Memory Usage Breakdown.
We take 8 experts, top-4 routing, and batch size 40 as an example to break down memory usage, and visualize the results in Figure LABEL:fig:mem_breakdown. We can draw some insights from it:
  1. $\star$
    The employment of pipeline-shared cache slightly increases memory consumption, while expert-specific fused kernel has no impact.
  2. $\star$
    Pipeline-shared cache can effectively reduce memory consumption under a data-centric setting since the memory usage would surpass Tutel without it.
  3. $\star$
    Our HEXA-MoE can reduce memory usage compared to baseline even without memory optimization, and optimizing it can further reduce memory consumption.
Latency Breakdown.
We set 4 experts, top-4 routing, and batch size 80 to break down training latency, and visualize results in Figure LABEL:fig:latency_breakdown. We can draw some insights as:
  1. $\star$
    The employment of expert-specific fused kernel, data-centric setting, and communication-computation overlap can effectively reduce latency. Meanwhile, the combination of data-centric setting and communication-computation overlap can further empower our HEXA-MoE with notably reduced latency.
  2. $\star$
    Expert-specific operators play a significant part in speeding up MoE computing since the latency for Tutel is much longer than our model-centric setting.
  3. $\star$
    Although the employment of pipeline-shared cache and memory optimization would slightly increase latency, they can also lead to significant memory reduction, as demonstrated in memory usage breakdown.
Figure 24: Memory and latency breakdown. We analyze the effectiveness of our proposed pipeline-shared cache, fused kernel, and memory-latency trade-off. We take Tutel as a baseline and visualize the differences with it.
6 Conclusion
In this paper, we analyze the limitations of current MoE libraries in detail and propose novel expert-specific operators as an alternative to GeMM so that MoE computing can be performed in an in-place manner with zero computation redundancy. We reformulate the forward and backward propagation of the MoE layer with these operators and replace conventional expert parallelism with tensor parallelism, i.e., splitting the parameters of the MoE layer along the hidden size dimension to distribute them on different devices. We provide GPU implementations for these operators in a highly parallelized manner, which can both save memory consumption and reduce overall latency. For different workload scales, we implement the MoE model under both data- and model-centric configurations. To tackle the teaser of heavy memory consumption in previous data-centric works, we propose a pipeline-shared cache for better memory efficiency. Homogeneous experiments show that HEXA-MoE can save $10\%$-$48\%$ memory usage while achieving $0.5$-$4.3$$\times$ speed up compared to state-of-the-art MoE libraries. To adapt HEXA-MoE to heterogeneous devices, we propose to assign different batch sizes or sub-dimensions of hidden sizes for different devices based on the heterogeneous computing capacity. Experiments show that our adaption can substantially minimize average latency and better utilize the whole computing capacity. Meanwhile, our HEXA-MoE empowers MoE computing to be conducted in an in-place manner, i.e., the routing choice would yield almost no impact on hardware workload, which can facilitate further research on MoE algorithms.
Appendix A Algorithm Details
We present the algorithm details for re-index vector construction as well as the expert-specific operators, and taking top-1 routing as an example for illustration, shown in Algorithm 1, 3, 2 and 4. In Algorithm 1, we re-arrange the routing choice vector $\mathcal{R}$ into re-indexed token vector $v$, along with the token index starting vector $idx$, which satisfies $idx[0]=0$ and $idx[E]=N^{\prime}$. We provide the length and range for the vectors in Table 4.
Table 4: Explanation for the auxiliary vectors.
vector length range
$\mathcal{R}$ $N$ $0\leq\mathcal{R}[i]<E$ for $0\leq i<N$
$v$ $N^{\prime}$ $0\leq v[i]<N$ for $0\leq i<N^{\prime}$
$idx$ $1+E$ $0\leq idx[i]<N^{\prime}$ for $0\leq i<1+E$
In Algorithm 3, 2 and 4, we assume that the feature dimension for each tensor is all divisible by BLK. Notice that when accessing the re-index vector $v$, we may get value -1, since the workload for each expert is dynamic. In this case, we would skip this index, and remain zero for the temporary loading variable.
We also provide the formulation for top-$k$ routing with our expert-specific operators, and take a comparison with top-$1$ routing. We denote the routing choice for top-$k$ as $\{\mathcal{R}_{i}(\bm{x})\}_{i=0}^{k-1}$, and other symbols keep consistent with Figure 5. We present the comparison on formulations for MoE forward and backward propagation in Table 5.
Table 5: Comparison between top-$1$ and top-$k$ routing for the formulation of MoE forward and backward with our expert-specific operators.
Stage Notation Layer Expert-Specific Formulation (top-$1$ routing) Expert-Specific Formulation (top-$k$ routing)
Forward 1 1st MLP $\bm{y}_{1}=\textit{ESMM}(\bm{x},\bm{W}_{1},\bm{b}_{1},\mathcal{R}(\bm{x}))$$\{\bm{y}_{1}^{i}\}_{i=0}^{k-1}:\bm{y}_{1}^{i}=\textit{ESMM}(\bm{x},\bm{W}_{1},\bm{b}_{1},\mathcal{R}_{i}(\bm{x}))$
2 Activation $\bm{y}_{2}=\mathcal{F}(\bm{y}_{1})$$\{\bm{y}_{2}^{i}\}_{i=0}^{k-1}:\bm{y}_{2}^{i}=\mathcal{F}(\bm{y}_{1}^{i})$
3 2nd MLP $\bm{y}=\textit{ESMM}(\bm{y}_{2},\bm{W}_{2},\bm{b}_{2},\mathcal{R}(\bm{x}))$ $\bm{y}=\sum_{i=0}^{k-1}\textit{ESMM}(\bm{y}_{2}^{i},\bm{W}_{2},\bm{b}_{2},\mathcal{R}_{i}(\bm{x}))$
Backward 4 2nd MLP $\frac{\partial\ell}{\partial\bm{b}_{2}}=\textit{ESS}(\frac{\partial\ell}{\partial\bm{y}},\mathcal{R}(\bm{x}))$ $\frac{\partial\ell}{\partial\bm{b}_{2}}=\sum_{i=0}^{k-1}\textit{ESS}(\frac{\partial\ell}{\partial\bm{y}},\mathcal{R}_{i}(\bm{x}))$
5 $\frac{\partial\ell}{\partial\bm{W}_{1}}=\textit{ESTMM}(\bm{x},\frac{\partial\ell}{\partial\bm{y}_{1}},\mathcal{R}(\bm{x}))$ $\frac{\partial\ell}{\partial\bm{W}_{1}}=\sum_{i=0}^{k-1}\textit{ESTMM}(\bm{x},\frac{\partial\ell}{\partial\bm{y}_{1}^{i}},\mathcal{R}_{i}(\bm{x}))$
6 $\frac{\partial\ell}{\partial\bm{y}_{2}}=\textit{ESMM}(\frac{\partial\ell}{\partial\bm{y}},\bm{W}_{2}^{T},\textit{null},\mathcal{R}(\bm{x}))$$\Big{\{}\frac{\partial\ell}{\partial\bm{y}_{2}^{i}}\Big{\}}_{i=0}^{k-1}:\frac{\partial\ell}{\partial\bm{y}_{2}^{i}}=\textit{ESMM}(\frac{\partial\ell}{\partial\bm{y}},\bm{W}_{2}^{T},\textit{null},\mathcal{R}_{i}(\bm{x}))$
7 Activation $\frac{\partial\ell}{\partial\bm{y}_{1}}=\frac{\partial\ell}{\partial\bm{y}_{2}}\odot\mathcal{F}^{\prime}(\bm{y}_{1})$$\Big{\{}\frac{\partial\ell}{\partial\bm{y}_{1}^{i}}\Big{\}}_{i=0}^{k-1}:\frac{\partial\ell}{\partial\bm{y}_{1}^{i}}=\frac{\partial\ell}{\partial\bm{y}_{2}^{i}}\odot\mathcal{F}^{\prime}(\bm{y}_{1}^{i})$
8 1st MLP $\frac{\partial\ell}{\bm{b}_{1}}=\textit{ESS}(\frac{\partial\ell}{\partial\bm{y}_{1}},\mathcal{R}(\bm{x}))$ $\frac{\partial\ell}{\bm{b}_{1}}=\sum_{i=0}^{k-1}\textit{ESS}(\frac{\partial\ell}{\partial\bm{y}_{1}^{i}},\mathcal{R}_{i}(\bm{x}))$
9 $\frac{\partial\ell}{\partial\bm{W}_{1}}=\textit{ESTMM}(\bm{x},\frac{\partial\ell}{\partial\bm{y}_{1}},\mathcal{R}(\bm{x}))$ $\frac{\partial\ell}{\partial\bm{W}_{1}}=\sum_{i=0}^{k-1}\textit{ESTMM}(\bm{x},\frac{\partial\ell}{\partial\bm{y}_{1}^{i}},\mathcal{R}_{i}(\bm{x}))$
10 $\frac{\partial\ell}{\partial\bm{x}}=\textit{ESMM}(\frac{\partial\ell}{\partial\bm{y}_{1}},\bm{W}_{1}^{T},\textit{null},\mathcal{R}(\bm{x}))$ $\frac{\partial\ell}{\partial\bm{x}}=\sum_{i=0}^{k-1}\textit{ESMM}(\frac{\partial\ell}{\partial\bm{y}_{1}^{i}},\bm{W}_{1}^{T},\textit{null},\mathcal{R}_{i}(\bm{x}))$
Algorithm 1 Constructing re-index vector
[…]
  Parallel for $i=0$ to $N-1$ do
    atomicAdd($ctr[\mathcal{R}[i]]$, 1)
  end for
  Parallel for $i=0$ to $E-1$ do
    $ctr[i]=\texttt{BLK}\cdot\lfloor ctr[i]\ /\ \texttt{BLK}\rfloor$
  end for $N^{\prime}=\sum\limits_{i=0}^{E-1}ctr[i]$
[…]
  Parallel for $i=0$ to $E-1$ do
    $ctr[i]=\texttt{BLK}\cdot\lfloor ctr[i]\ /\ \texttt{BLK}\rfloor$
  end for
  for $i=1$ to $E$ do
    $idx[i]=idx[i-1]+ctr[i-1]$
  end for
  Copy: $idx\_=idx$
  Parallel for $i=0$ to $N-1$ do
    $pos=\texttt{atomicAdd}(idx[\mathcal{R}[i]],1)$
    $v[pos]=i$
  end for
  Output: Tensor $v$ and $idx\_$
Algorithm 2 Expert-specific summation
[…]
  Initialize: Output tokens $y$ with shape $(E,D)$ initialized with 0.
  Parallel for $i=0$ to $E-1$ do
   Parallel for $j$ in range $(0,D,\texttt{BLK})\ \textbf{do}$
     $exp=\mathcal{R}[v[idx[i]]]$
    Initialize zero tensor $c$ with shape $(1,\texttt{BLK})$
[…]
    for $k$ in range $(idx[i],idx[i+1],\texttt{BLK})\ \textbf{do}$
     Parallel for $t=0$ to BLK  do
      load $x_{sub}[t]=x[v[k+t],j:j+\texttt{BLK}]$
     end for
      $c=c+\sum\limits_{t=0}^{\texttt{BLK}-1}x_{sub}[t]$
    end for
   Write back: $y[exp,j:j+\texttt{BLK}]=c$
   end for
  end for
  Output: Tensor $y$
Algorithm 3 Expert-specific matrix multiplication
[…]
  Initialize: Output tokens $y$ with shape $(N,D_{2})$ initialized with 0.
  Parallel for $i$ in range $(0,N^{\prime},\texttt{BLK})\ \textbf{do}$
   Parallel for $j$ in range $(0,D_{2},\texttt{BLK})\ \textbf{do}$
     $exp=\mathcal{R}[v[i]]$
[…]
    load $b_{sub}=b[exp,j:j+\texttt{BLK}]$
[…]
    for $k$ in range $(0,D_{1},\texttt{BLK})\ \textbf{do}$
[…]
     Parallel for $t=0$ to BLK  do
      load $x_{sub}[t]=x[v[i+t],k:k+\texttt{BLK}]$
     end for
     load $w_{sub}=w[exp,k:k+\texttt{BLK},j:j+\texttt{BLK}]$
      $c=c+x_{sub}\cdot w_{sub}$
    end for
    Parallel for $t=0$ to BLK  do
     Write back: $y[v[i+t],j:j+\texttt{BLK}]=c[t]$
    end for
   end for
  end for
  Output: Tensor $y$
Algorithm 4 Expert-specific transposed matrix     multiplication
[…]
  Initialize: Output $y$ with shape $(E,D_{1},D_{2})$ initialized with 0.
  Parallel for $i=0$ to $E-1$ do
   Parallel for $m$ in range $(0,D_{1},\texttt{BLK})\ \textbf{do}$
    Parallel for $n$ in range $(0,D_{2},\texttt{BLK})\ \textbf{do}$
      $exp=\mathcal{R}[v[idx[i]]]$
[…]
     for $k$ in range $(idx[i],idx[i+1],\texttt{BLK})\ \textbf{do}$
      Parallel for $t=0$ to BLK  do
       load $x_{sub}^{1}[t]=x_{1}[v[k+t],m:m+\texttt{BLK}]$
       load $x_{sub}^{2}[t]=x_{2}[v[k+t],n:n+\texttt{BLK}]$
      end for
       $c=c+x_{sub}^{1}.\text{transpose()}\cdot x_{sub}^{2}$
     end for
     Write back in parallel:
       $y[exp,m:m+\texttt{BLK},n:n+\texttt{BLK}]=c$
    end for
   end for
  end for
  Output: Tensor $y$
Appendix B Experimental Details
We provide the details of our CUDA program via enumerating the shape of the thread block and thread grid for expert-specific operators in a single MoE layer in Table 6.
Table 6: Shape of the thread block and thread grid for expert-specific operators in one MoE layer. We take top-$1$ routing as an example, where $N^{\prime}$ denotes the length of the re-index vector, which is slightly larger than $N$ and is divisible by BLK. Thread blocks are all defined with the same shape to facilitate the fused kernel.
Operator Input Output Thread Block Thread Grid
forward ESMM $\bm{x}$ $(N,D_{1})$ $\bm{y}$ $(N,D_{2})$ (WARP, TIMES) ($\lceil N^{\prime}/\texttt{BLK}\rceil$, $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
$\bm{w}$ $(E,D_{1},D_{2})$
$\bm{b}$ $(E,D_{2})$
backward ESMM $\bm{x}$ $(N,D_{2})$ $\bm{y}$ $(N,D_{1})$ (WARP, TIMES) ($\lceil N^{\prime}/\texttt{BLK}\rceil$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
$\bm{w}$ $(E,D_{2},D_{1})$
ESS $\bm{x}$ $(N,D_{2})$ $\bm{y}$ $(E,D_{2})$ (WARP, TIMES) ($E$, $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
ESTMM $\bm{x}_{1}$ $(N,D_{1})$ $\bm{w}$ $(E,D_{1},D_{2})$ (WARP, TIMES) ($\lceil D_{2}/\texttt{BLK}\rceil$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$, $E$)
$\bm{x}_{2}$ $(N,D_{2})$
ESFK $\bm{x}_{1}$ $(N,D_{1})$ $\bm{y}_{1}$ $(N,D_{1})$ (WARP, TIMES) ($E$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$, $\lceil N^{\prime}/\texttt{BLK}\rceil+\lceil D_{2}/\texttt{BLK}\rceil+$ $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
$\bm{x}_{2}$ $(N,D_{2})$ $\bm{y}_{2}$ $(E,D_{2})$
$\bm{w}_{1}$ $(E,D_{2},D_{1})$ $\bm{w}_{2}$ $(E,D_{1},D_{2})$
We also provide the Python code for the proxy task we used to examine the computing capacity of the heterogeneous devices. We adopt a for loop composed of large matrix multiplications with the same scale as the test program, as provided below:
import torch
import time
device = ’cuda’
size = 2048
times = 1024
start_time = time.time()
for j in range(times):
mat1 = torch.randn(size, size, device=device)
mat2 = torch.randn(size, size, device=device)
y = torch.matmul(mat1, mat2)
end_time = time.time()
print(end_time - start_time)
Appendix C Experimental Results
We provide the exact values for both memory usage analysis and average latency analysis in Table 7 and 8, respectively. Specifically, for latency analysis, we provide the average value for each case with 0.5k, 1k, 1.5k and 2k total steps.
We also visualize the validation accuracy of our method and Tutel under the same training time in Figure 25. We take Swin-MoE-Base with 4 experts and top-$1$ routing as the backbone network, training it on the homogeneous machine with the same batch size using Imagenet-1k dataset Deng et al. (2009). Since our HEXA-MoE can accelerate training process, its validation accuracy is much higher than baseline.
Table 7: Memory analysis with Tutel, MegaBlocks and HEXA-MoE on Swin-Transformer-MoE benchmark (Base and Small). Experiments are conducted on 2 homogeneous GPUs with automatic mixed precision in PyTorch and batch size 40 for all the experiments. We set the number of global experts to 8, and record the average GPU memory consumption (GB) on each device.
Method top-1 top-2 top-3 top-4 top-5 top-6 top-7 top-8
Base Tutel 12.7 13.9 15.3 16.0 17.4 19.0 20.3 21.8
MegaBlocks (MoE) 13.1 13.8 14.8 15.4 16.0 16.9 17.8 18.7
MegaBlocks (dMoE) 12.9 13.7 14.9 15.6 16.7 17.8 18.6 19.7
Ours (data-centric) 10.9 11.2 11.3 11.7 12.0 12.0 12.3 12.4
Ours (model-centric) 10.0 10.2 10.3 10.6 10.5 10.7 10.9 11.4
Small Tutel 9.0 10.0 11.0 11.6 12.7 13.8 15.0 16.0
MegaBlocks (MoE) 9.2 9.8 10.2 10.8 11.2 11.8 12.5 13.0
MegaBlocks (dMoE) 9.0 9.7 10.4 11.4 12.0 12.4 13.3 13.9
Ours (data-centric) 8.1 8.3 8.2 8.5 8.6 8.7 9.0 9.2
Ours (model-centric) 7.7 7.8 7.8 8.0 7.9 8.2 8.3 8.5
Table 8: Latency analysis for Tutel, MegaBlocks and HEXA-MoE on Swin-Transformer-MoE benchmark with base and small scale. Experiments are conducted on 4 homogeneous GPUs with 4 experts. We set different batch size (bs) for different models under different routing strategy to maximize the utilization of GPU memory. We record the average latency of one step (s) during training.
Method 0.5k 1k 1.5k 2k 0.5k 1k 1.5k 2k
Base top-1,bs=110 Tutel 2.96 2.90 2.47 2.40 top-2,bs=100 2.14 2.59 2.73 2.66
MegaBlocks (MoE) 2.10 2.66 2.57 2.43 2.40 2.58 2.51 2.49
MegaBlocks (dMoE) 2.06 2.02 2.13 2.19 2.47 2.72 2.63 2.55
Ours (model-centric) 1.52 1.51 1.51 1.51 1.61 1.60 1.60 1.60
Ours (data-centric) 1.01 0.99 0.99 0.99 1.17 1.16 1.16 1.16
top-3,bs=90 Tutel 2.23 2.43 2.48 2.54 top-4,bs=80 2.27 2.34 2.20 2.18
MegaBlocks (MoE) 2.09 2.31 2.35 2.17 2.11 2.18 2.05 2.03
MegaBlocks (dMoE) 2.55 2.53 2.44 2.41 2.00 2.19 2.23 2.12
Ours (model-centric) 1.70 1.69 1.69 1.69 1.83 1.82 1.82 1.82
Ours (data-centric) 1.32 1.31 1.31 1.30 1.42 1.41 1.41 1.41
Small top-1,bs=140 Tutel 3.59 3.63 3.62 3.01 top-2,bs=130 2.83 2.96 2.72 2.53
MegaBlocks (MoE) 3.41 3.64 3.72 3.23 3.04 3.26 3.11 3.02
MegaBlocks (dMoE) 3.51 3.58 3.76 3.47 2.25 3.10 3.01 2.87
Ours (model-centric) 1.34 1.34 1.33 1.33 1.49 1.49 1.48 1.48
Ours (data-centric) 0.69 0.68 0.68 0.68 0.87 0.86 0.86 0.85
top-3,bs=120 Tutel 2.23 2.92 3.17 3.28 top-4,bs=110 3.09 2.86 3.03 3.00
MegaBlocks (MoE) 2.63 2.79 2.92 3.07 3.04 3.09 3.01 2.91
MegaBlocks (dMoE) 3.08 3.24 3.38 3.09 3.03 3.19 3.18 2.91
Ours (model-centric) 1.61 1.60 1.59 1.59 1.67 1.67 1.68 1.68
Ours (data-centric) 1.07 1.05 1.05 1.05 1.09 1.08 1.07 1.07
Figure 25: Comparison on validation accuracy for Tutel and HEXA-MoE with the same training time. HEXA-MoE has a significant advantage over Tutel due to the highly improved efficiency on computation.

### v4

@@
-
+Hexa-MoE: Efficient and Heterogeneous-aware Training for Mixture-of-Experts
 Abstract
-Mixture-of-Experts (MoE) has emerged as a practical approach to scale up parameters for the Transformer model to achieve better generalization while maintaining a sub-linear increase in computation overhead.
-Current MoE models are mainly built with expert parallelism on distributed devices.
-However, it usually depends on homogeneous devices to deploy and suffers from heavy communication overhead and computation redundancy.
-In this paper, we explore developing a Heterogeneous-aware EXpert Allocation framework, HEXA-MoE, with significantly enhanced computing efficiency.
+Mixture-of-experts (MoE) has emerged as a practical approach to scale up parameters for Transformer model to achieve better generalization while keeping computational efficiency.
+Current MoE models are mainly deployed with expert parallelism on distributed devices, which typically depends on homogeneous devices, and suffers from heavy communication overhead as well as computation redundancy under scaled workload.
+To tackle these teasers, we propose a Heterogeneous-aware EXpert Allocation framework, HEXA-MoE, with significantly enhanced efficiency.
 It contains two components: ($1$) Expert-Specific Operators.
-We replace the typical general matrix multiplication or grouped matrix multiplication interfaces with our operators, which allows the computing to be performed in an in-place manner with ZERO redundancy.
+We replace the typical GeMM or grouped GeMM interfaces with our proposed expert-specific operators, making expert computing to be performed in an in-place manner with almost no redundant FLOPs.
 ($2$) Adaptive Data- and Model-Centric configurations for different workload scales.
-Specifically, we introduce a pipeline-shared cache on each device to tackle the heavy memory consumption in the existing data-centric MoE library.
-Comprehensive experiments on the Swin-MoE benchmark consistently reveal the effectiveness of our HEXA-MoE framework, i.e., reducing $10\%\sim 48\%$ memory consumption and achieving 0.54.3×0.5$\sim$ 4.3$\times$0.5 ∼ 4.3 × speed up compared to current state-of-the-art MoE libraries.
-Furthermore, we examine our HEXA-MoE with heterogeneous devices for both data- and model-centric settings.
-Promising results show that employing optimal parallel configuration with HEXA-MoE on heterogeneous devices can substantially minimize overall latency.
+We introduce a memory-efficient computation-communication overlapping scheme to tackle the heavy memory consumption in current data-centric libraries, which can accelerate training with heavy workloads.
+Comprehensive experiments on the Swin-MoE benchmark consistently reveal the effectiveness of HEXA-MoE, i.e., reducing $10\%\sim 48\%$ memory consumption and achieving 0.54.3×0.5$\sim$ 4.3$\times$0.5 ∼ 4.3 × speed up compared to current state-of-the-art MoE libraries.
+We further examine HEXA-MoE on heterogeneous devices, and promising results show that employing optimal parallel configuration can better utilize global computation resources, and substantially minimize overall latency.
 Codes are available at here.
-footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
-Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
+Machine Learning, ICML
 1 Introduction
-Transformers have become the de facto architecture for a vast range of machine learning tasks including natural language process Vaswani (2017); Raffel et al.
-(2020), computer vision Liu et al.
-(2021); He et al.
-(2022b) and multi-modal learning Li et al.
-(2022); Liu et al.
-(2024).
-To scale up model parameters for stronger learning capacity and better generalization following the scaling law Kaplan et al.
-(2020), Mixture-of-Experts (MoE) Jiang et al.
-(2024) has been demonstrated as a practical and computation-efficient approach.
-For a single MoE layer, the feed-forward network (FFN) is expanded to a group of experts compared to the dense counterpart, activated selectively during computing.
+Figure 1: Convergence Analysis. HEXA-MoE can significantly surpass Tutel on MoE training due to the specialized designs.
+Transformers have become the de facto architecture for a vast range of machine learning tasks including natural language process (Vaswani, 2017; Raffel et al., 2020), computer vision (Liu et al., 2021; He et al., 2022b) and multi-modal learning (Li et al., 2022; Liu et al., 2024).
+To scale up model parameters for stronger learning capacity and better generalization following the empirical scaling law (Kaplan et al., 2020), Mixture-of-experts (MoE) (Jiang et al., 2024) has been demonstrated as a practical and computation-efficient approach.
+For a single Transformer layer, MoE expands the feed-forward network (FFN) to a group of experts with the same architecture and different parameters, activated selectively during computing.
 This dynamic nature poses peculiar challenges to system design.
@@
   1.
-Distributed computing is required for MoE due to its sheer size, and expert parallelism Lepikhin et al.
-(2020) is the most commonly used technique, which distributes the experts in one layer on different devices.
+Distributed computing is required due to its sheer size, and expert parallelism (Lepikhin et al., 2020) is the most commonly used technique, which distributes the experts in one layer on different devices.
   2.
-Since the workload varies dynamically for each expert in each layer, we have to re-organize the local data batches for expert computing via dispatch and combine operations.
-Specifically, we have to calibrate the workload for each expert to utilize the general matrix multiplication (GeMM) or grouped GeMM interface.
-  3.
-Since each token may activate any one or several experts and different experts may be distributed on different devices in expert parallelism, it is necessary to communicate among devices before and after each MoE layer via all-to-all communication.
-Figure 2: Tensor Parallelism vs Expert Parallelism for MoE layer.
-We take $4$ experts on $2$ devices as an example and visualize the parameter distribution of $2$ MLPs.
-For expert parallelism, each device can only keep several experts, while for tensor parallelism the parameter allocation is more flexible since the partition of parameters is along the hidden dimension.
-These peculiar designs lead to inefficiency for both training and inference in MoE computing.
-On the one hand, dispatch and combine operations result in unnecessary token padding or discarding Gale et al.
-(2023), leading to both redundant computation and compromise on model performance.
-Meanwhile, conventional expert parallelism depends on all-to-all communication, which requires synchronization and occupies a lot of latency Li et al.
-(2023).
-On the other hand, current expert parallelism for MoE models is mainly deployed on homogeneous devices.
-However, a cluster equipped with the latest homogeneous devices is usually much more expensive than heterogeneous devices due to the iteration of GPU development and the decrease on prices for old-version GPUs.
-If we want to adapt expert parallelism to heterogeneous devices, it would be necessary to rearrange the expert distribution to utilize the total computing capacity better.
-However, the inherent dynamic property of the MoE layer makes it difficult to implement, as the workload of each expert changes dynamically in each step during computing He et al.
-(2022a); Li et al.
-(2023).
-Meanwhile, since heterogeneous devices are cheaper and easier to access if MoE computing can be re-designed to be heterogeneous-aware, it would benefit wider deployment.
-This paper proposes HEXA-MoE, a heterogeneous-aware framework for efficient MoE computing.
-We first find that if compiling the MoE layer into GeMM like other operators such as Linear, Convolution, etc., the unnecessary components such as token padding or discarding are unavoidable.
-To tackle it, we investigate the forward and backward propagation for the MoE layer thoroughly, and propose to replace current GeMM or grouped GeMM Gale et al.
-(2023) interfaces with expert-specific operators.
-Specifically, all the computations related to the MoE layer can be summarized as built on 3 operators, i.e., expert-specific matrix multiplication (ESMM), expert-specific summation (ESS) and expert-specific transposed matrix multiplication (ESTMM).
-The forward propagation only requires ESMM, while backward requires all the $3$ operators.
-We implement them in a highly-parallelized manner on GPU, and distribute the MoE layer with tensor parallelism Narayanan et al.
-(2021) as an alternative to expert parallelism, shown in Figure 2.
-HEXA-MoE also considers different workload scales for MoE computing and is adapted to both data- and model-centric configurations.
-Specifically, the difference lies in the content of communication.
-For model-centric, local data batches are synchronized among devices, and model parameters are kept still, while for data-centric, local parameters are gathered to each device, while local data batches are kept still Liu et al.
-(2023).
-When the scale of data is larger than the model parameters, the data-centric setting would perform better.
-However, previous work preserves all the gathered MoE parameters on each device for backward propagation, leading to huge memory consumption.
-To achieve better memory efficiency while maintaining high computing speed, we propose to introduce a pipeline-shared cache on each device, which is updated dynamically during both forward and backward propagation, and the all-gather communication can be overlapped with non-MoE operations such as attention.
-Essentially, a data-centric setting can be viewed as data parallelism, which is easy to adapt to heterogeneous devices.
-A model-centric setting is employed when the size of model parameters is larger than the data workload.
-In this case, all gathered communication is applied to the data batches, and tensor parallelism is applied to each MoE layer.
-Each device processes all the data batches using the parameter chunk it preserves.
-This can also be employed in heterogeneous devices via adaptively allocating expert parameters, as illustrated in Figure 2.
+The dynamic workload for each expert makes it necessary to employ additional dispatch and combine operations in expert parallel to utilize the general matrix multiplication (GeMM) or grouped GeMM interface.
+  3. Dispatch and combine rely on synchronous all-to-all communication to assign tokens to distributed experts.
+These peculiar designs lead to apparent inefficiency for MoE training.
+On the one hand, current MoE libraries are either static like Tutel (Hwang et al., 2023), or dynamic like MegaBlocks (Gale et al., 2023).
+The former depends on an additional hyper-parameter named expert capacity to calibrate the workload for different experts, which suffers from redundant memory allocation or access, while the latter tends to raise out of memory (OOM) error during runtime due to its uncertainty.
+Meanwhile, the synchronous all-to-all communication can occupy over $40\%$ runtime with scaled model and devices (Hwang et al., 2023).
+On the other hand, expert parallelism is mainly deployed on homogeneous devices.
+However, a homogeneous cluster with cutting-edge devices is much more expensive than heterogeneous ones due to the fast iteration of modern GPU and the decrease on prices for outdated ones.
+Adapting expert parallelism to heterogeneous devices would rely on rescheduling expert placement to utilize global computing resources sufficiently.
+However, the inherent dynamic property of MoE makes it difficult to implement, as the workload of each expert changes dynamically in each step (He et al., 2022a; Li et al., 2023).
+Since heterogeneous devices are cheaper and easier to access, if MoE training can be refactored to be heterogeneous-aware, it would advance wider deployment.
+We propose HEXA-MoE, a completely-static MoE training framework with minimized memory consumption and heterogeneous-awareness.
+We first find that if implementing MoE layer with GeMM, the unnecessary components such as token padding or discarding are unavoidable, while grouped GeMM (Gale et al., 2023) would lead to dynamic runtime behavior and uncertainty.
+To tackle it, we investigate the forward and backward propagation for the MoE layer thoroughly, and propose to replace (grouped) GeMM with expert-specific operators.
+Specifically, the basic MoE operations are essentially build upon $3$ operators, i.e., expert-specific matrix multiplication (ESMM), summation (ESS) and transposed matrix multiplication (ESTMM).
+The forward propagation only requires ESMM, while backward requires all.
+We implement them with specialized GPU kernels, and distribute the MoE layer with tensor parallelism (Narayanan et al., 2021) as an alternative to expert parallelism, which is easy to be heterogeneous-aware.
+HEXA-MoE also considers different workload scales for MoE training, and is adapted to both data- and model-centric configurations.
+Their difference lies in the content of communication.
+For model-centric, local mini-batches are synchronized among devices and model parameters are kept still, while for data-centric, local parameters are gathered for each device, while local mini-batches are kept still (Liu et al., 2023).
+Data-centric setting outperforms model-centric with scaled workloads.
+However, previous library preserves all the gathered parameter shards on each device for backward pass, leading to huge memory consumption.
+To tackle this teaser, we design a novel memory-efficient communication-communication overlapping scheme utilizing a pipeline-shared cache on each device, which is updated dynamically for both forward and backward pass.
+Both settings are completely static in runtime, and the workload of each device can be easily determined from batch size or tensor parallelism configuration (sub-dimension of FFN intermediate size), making it easy to be adapted to heterogeneous devices via specialized expert allocation.
 Our contributions can be summarized as follows:
   1. $\star$
-    We propose to refactor MoE computing with expert-specific operators, enabling it to be performed in an in-place manner without the requirement of dispatch and combine operations.
-In this way, ZERO computation redundancy would be introduced.
-We implement these operators in a highly parallelized manner on GPU.
+    A completely static MoE training framework with minimized memory footprint.
+To our best knowledge, HEXA-MoE is the first MoE library with completely static and minimal memory footprint, enabling in-place computing with specialized kernel, and employing tensor parallelism for distributed training, delivering exact result with faster speed, as shown in Figure 1.
   2. $\star$
-    We develop HEXA-MoE based on the expert-specific operators for both data- and model-centric settings, considering different scales of workload, and propose to employ pipeline-shared cache for data-centric setting to improve memory efficiency.
-Experiments with Swin-MoE show that HEXA-MoE can reduce $10\%\sim 48\%$ memory usage and achieve 0.54.3×0.5$\sim$ 4.3$\times$0.5 ∼ 4.3 × speed up compared to state-of-the-art MoE computing libraries.
+    Advanced efficiency.
+HEXA-MoE is built upon our proposed expert-specific operators, which tackles scaled workload with data-centric strategy via memory-efficient communication-computation overlapping.
+Experiments on Swin-MoE show that HEXA-MoE can reduce $10\%\sim 48\%$ memory usage and achieve 0.54.3×0.5$\sim$ 4.3$\times$0.5 ∼ 4.3 × speed up compared to SOTA MoE libraries.
   3. $\star$
-    Our HEXA-MoE transforms conventional expert parallelism into pure data or tensor parallelism, making it heterogeneous-aware.
-Experiments show that our HEXA-MoE can be effectively adapted to heterogeneous devices, and employing the optimal parallel configuration can substantially minimize overall latency.
+    Heterogeneous-awareness.
+HEXA-MoE transforms conventional expert parallelism into data or tensor parallelism, making it heterogeneous-aware.
+Experiments show that HEXA-MoE can be effectively adapted to heterogeneous devices and substantially minimize overall latency via employing optimal parallel configuration.
+2 Related Works
+Figure 2: Comparison between conventional and expert-specific formulation for MoE computing.
+We take top-1 routing for illustration and present the corresponding relation of each formula in the MoE forward and backward propagation.
+Scaling Transformers with MoE.
+Transformer models can be scaled up via the Mixture-of-Experts (MoE) for better learning capacity and generalization while enjoying a sub-linear increase in computation overhead due to its sparsity.
+It has been proven in many research fields such as natural language processing (Fedus et al., 2022; Jiang et al., 2024; Du et al., 2022), machine vision (Riquelme et al., 2021; Fan et al., 2022) and multi-modal learning (Bao et al., 2022).
+Open-Source MoE Training Libraries.
+FastMoE (He et al., 2021) pioneered the first PyTorch open-source implementation for distributed MoE.
+Based on it, FasterMoE (He et al., 2022a) proposes dynamic shadowing and smart scheduling to tackle load imbalance and improve parallelism.
+Tutel (Hwang et al., 2023) implements switchable parallelism and dynamic pipelining, improving adaptability and scalability.
+MegaBlocks (Gale et al., 2023) proposes block-sparse operations and corresponding GPU kernels to mitigate dynamic routing issues in MoEs.
+Improving efficiency for MoE computing.
+MoE computing faces significant challenges in communication and memory.
+Recent research has explored alleviating them from various perspectives: Lina (Li et al., 2023) proposes to dynamically schedule resources to reduce latency and tackle the all-to-all bottleneck.
+Janus (Liu et al., 2023) proposes a data-centric paradigm that replaces traditional all-to-all communication with asynchronous expert fetching, enabling overlapped communication and computation for enhanced parallelism.
+ScheMoE (Shi et al., 2024) proposes a generic scheduling framework for optimal communication and computation scheduling during MoE training.
+MPMoE (Zhang et al., 2024) accelerates MoE training through adaptive and memory-efficient pipeline parallelism.
+SmartMoE (Zhai et al., 2023) proposes an efficient searching algorithm to identify optimization opportunities within an expanded hybrid parallelism space, tailored for data-sensitive MoE models.
+PipeMoE (Shi et al., 2023) adaptively pipelines communications and computations in MoE to mask communication latency.
+It also provides an optimal strategy to determine pipeline degree to minimize overall iteration time.
+Principles of GPU Acceleration
+Modern GPUs provide massive threads for parallel execution, grouped into thread-blocks, and executed on streaming multiprocessors (SMs).
+GPUs have a memory hierarchy, outlined as large but slow-accessed high bandwidth memory (HBM), and small but faster-accessed shared memory (SRAM).
+Matrix multiplication is optimized on GPU with tiling, i.e., partitioning the output matrix into small $2$D blocks, where each block is computed by a thread-block in parallel.
+The size of individual blocks can be adjusted to improve runtime performance.
+3 Method
+We first formulate the forward and backward propagation for a single MoE layer with GeMM and expert-specific operators respectively (Sec.
+3.1), which deliver exact results in different manner.
+After that, we expound the details for our specialized kernel (Sec.
+3.2), where a novel expert-specific fused operator is introduced for parallel MoE backward pass.
+These empower HEXA-MoE with high computational efficiency.
+Next, we consider different workload scales, and adapt HEXA-MoE efficiently to both data- and model-centric settings (Sec.
+3.3).
+Finally, we adapt HEXA-MoE to heterogeneous devices, and provide an expert allocation approach to minimize the average latency so as to utilize better heterogeneous computing capacities (Sec.
+3.4).
+3.1 MoE Computing Formulation
+MoE Computing with GeMM.
+To employ GeMM for expert computing, we need to calibrate the amount of dispatched tokens for each expert via token padding or discarding.
+Taking top-1 routing as an example, we formulate the forward and backward propagation in Figure 2, denoting $\ell$ as loss value and $\boldsymbol{x}^{e}$ as the $N_{\textrm{e}}$ tokens dispatched to expert $e$.
+In backward propagation, the gradients of the output have been provided by the auto-differentiation program.
+Based on dispatch and combine, all variables can be derived via basic matrix operations such as summation and multiplication.
+Although we can utilize the high-performance GeMM interface, these operations are memory inefficient, since the workload of each expert varies dynamically in each step, and token padding or discarding has to be employed to construct new mini-batches for local experts.
+It can be overcome by our expert-specific design.
+MoE Computing with Expert-Specific Operators.
+Based on the above formulations, we propose to refactor the MoE workflow in an in-place manner to address the teaser of inefficiency and dynamic.
+Specifically, we find that the forward and backward propagation of a single MoE layer can be reformulated with $3$ basic specialized operators from an expert-specific perspective, namely expert-specific operators.
+We take top-$1$ routing for illustration, while formulations for top-$k$ routing are provided in Appendix B.
+(a) Re-Index Vector Construction
+(b) Expert-Specific Matrix Multiplication
+(c) Expert-Specific Summation
+(d) Expert-Specific Transposed Matrix Multiplication
+Figure 3: Illustration of the proposed operators.
+We take $10$ tokens, $4$ global experts, and tiling size $4$ as an example.
+For ESTMM, the $2$ input batches are in a re-indexed format, while for others both the raw batch and re-index vector are provided.
+  1.
+Expert-specific matrix multiplication (ESMM): Given input $\boldsymbol{x}$ with $N$ tokens, routing choice $\mathcal{R}(\boldsymbol{x})$, weight $\boldsymbol{W}$ and bias $\boldsymbol{b}$, the output $\boldsymbol{y}=\textit{ESMM}(\boldsymbol{x},\boldsymbol{W},\boldsymbol{b},\mathcal{R}(\boldsymbol{x}))$, where $\boldsymbol{y}_{i}$ is derived from $\boldsymbol{x}_{i}$, $\boldsymbol{W}_{\mathcal{R}(\boldsymbol{x}_{i})}$ and $\boldsymbol{b}_{\mathcal{R}(\boldsymbol{x}_{i})}$.
+  2.
+Expert-specific summation (ESS): Given input $\boldsymbol{x}$ with $N$ tokens and routing choice $\mathcal{R}(\boldsymbol{x})$, the output $\boldsymbol{y}=\textit{ESS}(\boldsymbol{x},\mathcal{R}(\boldsymbol{x}))$, where tokens routed to expert $e$ are added up and recorded in $\boldsymbol{y}[e]$.
+  3.
+Expert-specific transposed matrix multiplication (ESTMM): the inputs include $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, both with $N$ tokens, and the routing choice $\mathcal{R}(\boldsymbol{x})$.
+Both $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$ are the ESMM result of $\boldsymbol{x}$, thus sharing the same routing choice.
+The output $\boldsymbol{y}=\textit{ESTMM}(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\mathcal{R}(\boldsymbol{x}))$.
+For the $i$-th channel of $\boldsymbol{x}_{1}$ and $j$-th channel of $\boldsymbol{x}_{2}$, we first prepare a zero vector $\boldsymbol{c}$ with $E$ elements, and accumulate $\boldsymbol{x}_{1}[m,i]\cdot\boldsymbol{x}_{2}[m,j]$ to $\boldsymbol{c}[\mathcal{R}(\boldsymbol{x}_{m})]$ for all $0\leq m<N$.
+After that we assign $\boldsymbol{c}[k]$ to $\boldsymbol{y}[k,i,j]$ for all $0\leq k<E$.
+Based on these expert-specific operators, we can now implement MoE computing in a novel in-place manner, as formulated in Figure 2.
+We compare each stage between our method and conventional formulation, and the output of each token is ensured to be consistent for both cases.
+In forward pass, ESMM serves as an alternative to GeMM, while in backward pass, the gradient of $\boldsymbol{y}$ is also provided by the auto-differentiation program.
+ESS performs as an alternative to tensor summation, while ESMM and ESTMM are designed as $2$ specialized cases for matrix multiplication from the expert-specific perspective.
+The expert-specific design enables HEXA-MoE to be a completely static framework at runtime, introducing almost no redundant FLOPs.
+3.2 Optimized CUDA Implementations
+Re-Index based Expert-Specific Operators.
+To implement tiled matrix multiplication for the expert-specific operators, the HBM I/O should be re-directed, since naively implement them cannot fully utilize GPU locality as well as tensor cores, which restricts runtime performance.
+In HEXA-MoE, we introduce re-index vector as I/O guidance, illustrated in Figure 3(a).
+Specifically, we re-organize the order of token indices based on the routing choice, i.e., gathering the token indices routed to the same expert into a sub-vector, and padding it with -1 to make it divisible by the tiling size.
+Algorithm details are provided in Appendix B.
+For ESMM, we illustrate it in Figure 3(b), where a thread-block first loads a sub-vector, followed by input tokens based on the vector values, as well as the corresponding expert parameters.
+Since the loaded tokens are routed to the same expert, we only need to load weights for one expert.
+We then accumulate the dot product results along the dimension of the input feature, and write the result back to HBM guided by the sub-vector.
+For ESS, we illustrate it in Figure 3(c), where each thread-block is assigned with certain channels of one expert, with tokens specialized by the sub-vectors for that expert.
+After accumulating all the assigned tokens, it writes the result back to HBM.
+For ESTMM, we illustrate it in Figure 3(d), where input batches are presented in a re-indexed manner.
+The two inputs here are the ESMM results of the same tensor, sharing the same re-index vector.
+Each thread block loads certain channels of both inputs for tokens routed to the same expert.
+The cumulative dot production result is write back to the corresponding area on HBM after computing.
+More algorithm details of this section are provided in Appendix B.
+Memory Optimization for Top-$k$ Routing.
+We find that when extending the routing policy from top-$1$ to top-$k$, the memory footprint would increase significantly.
+To tackle it, we only enlarge the memory allocation for the intermediate tokens to $k$ times.
+The input and output tensors (and their gradients) are the cumulation of $k$ ESMM results, which can either be implemented in serial, or employ the atomic add interface [^1] Some data types may not be supported for atomic_add on NVIDIA GPUs.
+We initialize the output tensor with float32, and transform it to the target type after computing., which lead to similar runtime.
+(a) Training pipeline under data-centric configuration.
+(b) Pipeline-shared cache before and after all gather.
+Figure 4: Visualization of the training pipeline and shared cache in data-centric setting.
+Each device copies the kept parameter shard to the cache before all gather communication, and after that it can access the whole parameters of an MoE layer.
+Fused Kernel in Backward Propagation.
+For a single layer MLP $\boldsymbol{y}=\boldsymbol{x}\cdot\boldsymbol{W}+\boldsymbol{b}$, the gradients for $\boldsymbol{x}$, $\boldsymbol{W}$ and $\boldsymbol{b}$ can be computed in parallel.
+Similarly, we can integrate ESS, ESTMM, and ESMM together into one kernel, namely expert-specific fused kernel (ESFK) for enhanced parallelism.
+However, directly integrating them is difficult, since the shape of thread-blocks and the implementation details for each operator vary a lot.
+To this end, we set the shape of thread-block for each operator to be the same, and expand one dimension for ESMM and ESS so that the thread grids are all $3$-dimensional.
+Taking a single FFN layer under top-$1$ routing as an example, where we present the shape of allocated thread block and thread grid for both forward and backward propagation in Table 6 in Appendix C.
+By shape transposing or dim expanding for individual thread-blocks, we can align the 1st and 2nd dim while aggregate the 3rd dim for an integrated thread-grid.
+In this way, the backward pass for an MoE layer can be implemented with only $2$ fused kernels and an element-wise dot production.
+3.3 Parallel Strategy for Different Workload Scale
+Data Parallelism for Data-Centric Setting.
+Data-centric is a practical and efficient approach for training deep neural network models with heavy workloads.
+HEXA-MoE distributes MoE parameters to different devices based on the partition of FFN intermediate size, namely tensor parallelism.
+In data-centric setting, each device gathers the whole MoE parameters of one layer from other devices, and computes locally, similar to data parallelism.
+Although data-centric MoE training has been explored by Janus (Liu et al., 2023), the teaser of memory efficiency has not been fully tackled.
+Janus pre-fetchs the required MoE parameters for each layer progressively in forward pass, and preserves all of them locally for backward pass so that no communication would be required in backward.
+However, the sheer size of MoE parameters can lead to a significant increase in memory usage, making it inefficient for deployment.
+To tackle this issue, we propose to allocate an additional region on HBM of each GPU to dynamically cache the gathered MoE shards for each pipeline stage, named as pipeline-shared cache.
+Since each device keeps a subset of the FFN intermediate size for all experts in each layer, we employ all gather communication among devices before computing for an MoE layer, preparing for the required full parameters in both forward and backward pass, as illustrated in Figure 4(a) and 4(b).
+For better parallelism, all gather communication can be overlapped with other operations such as attention and router, as shown in Figure 4(a).
+In this way, each device would not have to preserve the full MoE parameters for backward pass, while communication overhead can also be overlapped, therefore both memory efficiency and computing efficiency can be achieved.
+Tensor Parallelism for Model-Centric Setting.
+For small scale of workload less than model parameters, model-centric configuration turns out to be more efficient than data-centric.
+To distribute MoE parameters on multiple devices, we modify the classical tensor parallelism with our proposed expert-specific operators.
+All gather communication is employed to synchronize local data batches before and after each MoE layer.
+We also distribute each MoE layer among different devices along the FFN intermediate size for each expert, and during MoE computing, each device computes the all gathered data batches with only the local MoE parameter chunk using ESMM, after that the output tokens are all reduced with sum operation in forward pass.
+During backward propagation, all gather and all reduce communications are interchanged, while ESMM are replaced by ESMM, ESS, and ESTMM to get the gradients for input tokens, bias, and weights, respectively, which can also be replaced by the fused operator ESFK.
+Figure 5: Average memory usage for training Swin-Transformer-MoE models.
+We take 8 global experts and examine all cases from top-1 to top-8 routings.
+Experiments are conducted on 2 homogeneous GPUs using automatic mixed precision in PyTorch.
+The batch size is set to 40 for all cases.
+We record the average GPU memory consumption (GB) on each device.
+3.4 Heterogeneous-aware Expert Allocation
+Workload Division for different configurations.
+For data-centric, the workload mainly depends on local batch size, as it can be essentially viewed as data parallelism.
+For model-centric, the workload of each device mainly depends on the allocated sub-dimension of FFN intermediate size for each expert, as expert-specific operators enable the implementation of tensor parallelism.
+Practical Expert Allocation on Heterogeneous Devices.
+We propose a specialized expert allocation approach to utilize heterogeneous computing resources, by adjusting the workload of each device.
+Specifically, we have to first examine the computing capacity of each device by averaging its latency on a benchmark task with heavy computing, such as large matrix multiplication.
+Denoting the results as $\{t_{i}\}_{i=0}^{N-1}$ for $N$ GPUs.
+For data-centric, we assign different local batch size $\{B_{i}\}_{i=0}^{N-1}$ to different devices based on the examined latency, as illustrated in Equation 1:
+$B_{i}=\frac{1/t_{i}}{\sum_{j=0}^{N-1}1/t_{j}}\cdot B_{\textrm{global}},$ (1)
+where we denote $B_{\textrm{global}}$ as the global batch size, i.e., $B_{\textrm{global}}=\sum_{j=0}^{N-1}B_{i}$.
+For model-centric, we denote the sub-dimension of hidden size for one MoE layer on each device as $\{h_{i}\}_{i=0}^{N-1}$ with total hidden size $H$.
+The assigned sub-dimension $h_{i}$ on device $i$ can be derived as Equation 2.
+$h_{i}=\frac{1/t_{i}}{\sum_{j=0}^{N-1}1/t_{j}}\cdot H.$ (2)
+To ensure that both $\{B_{i}\}_{i=0}^{N-1}$ and $\{h_{i}\}_{i=0}^{N-1}$ are integers, we need to further round them up or down while making sure that the summation is exactly $B_{\textrm{global}}$ or $H$.
+Table 1: Details of the machines and GPUs used for both homogeneous and heterogeneous experiments.
+Notation Number & Hardware Specs Memory
+CPU $M_{\textrm{homo}}$ 2×2$\times$2 × Intel Xeon Platinum 8352V 2.10GHz 1008 GB
+$M_{\textrm{hete}}$ 2×2$\times$2 × Intel Xeon Gold 6130 2.10GHz 62.5 GB
+GPU $D$ (in $M_{\textrm{homo}}$) 4×4$\times$4 × NVIDIA GeForce RTX 4090 24 GB
+$D_{0}$ (in $M_{\textrm{hete}}$) 1×1$\times$1 × NVIDIA TITAN RTX 24 GB
+$D_{1}$ (in $M_{\textrm{hete}}$) 1×1$\times$1 × NVIDIA GeForce RTX 2080 Ti 11 GB
+4 Experiments
+4.1 Experimental Setup
+We implement all the experiments using 2 machines, namely $M_{\textrm{homo}}$ and $M_{\textrm{hete}}$.
+Specifically, $M_{\textrm{homo}}$ is composed of 4 homogeneous GPUs, while $M_{\textrm{hete}}$ contains 2 heterogeneous GPUs.
+Details for the machines and GPUs are provided in Table 1.
+We take the training process of Swin-Transformer-MoE as the benchmark for all the experiments, including memory analysis, latency analysis, heterogeneous computing analysis, and ablation studies.
+Apart from heterogeneous computing analysis, all the experiments are conducted on $M_{\textrm{homo}}$.
+We also adopt both the Small and Base scales for the Swin-MoE model, following Tutel (Hwang et al., 2023).
+Experiments are conducted with PyTorch, taking nccl as the communication backend.
+Batch size is recorded as the workload of single device, and automatic mixed precision is used for training.
+atomicAdd is employed for HEXA-MoE to aggregate the expert computing results for each token.
+Figure 6: Average latency for training Swin-Transformer-MoE models.
+Experiments are conducted on 4 homogeneous GPUs with 4 global experts.
+We set different batch sizes for different models under different routing strategies to maximize the utilization of GPU memory.
+We record the average latency for one step (s) during training with 2k steps in total.
+4.2 Experimental Results
+Memory Analysis on Homogeneous Devices.
+We analyze the memory footprint for different MoE libraries on homogeneous devices with 8 global experts, and examine from top-$1$ to top-$8$ routing.
+Results are visualized in Figure 5.
+HEXA-MoE can reduce $10\%$-$48\%$ memory footprint compared to Tutel (Hwang et al., 2023) and MegaBlocks (Gale et al., 2023), and the memory footprint for model-centric is slightly less than data-centric owing to the pipeline-shared cache.
+From top-$1$ to top-$k$ routing, the memory footprint increase of HEXA-MoE is more gentle than others, owing to the introduction of memory optimization.
+Memory saving mainly benefits from the in-place design of expert-specific operators and memory optimization.
+Latency Analysis on Homogeneous Devices.
+We analyze the average latency for one training step, and present the results in Figure 6.
+To fully utilize the GPU HBM, we set different batch sizes for different routing configurations and different model scales.
+HEXA-MoE achieves $0.5$-$4.3$$\times$ speed up compared to Tutel (Hwang et al., 2023) and MegaBlocks (Gale et al., 2023).
+Since the scale of workload is larger than MoE parameters in our experiments, the average latency under a data-centric setting is less than model-centric in these cases, and the advantage of HEXA-MoE turns more obvious while enlarging batch size.
+To better visualize the latency comparison between data-centric and model-centric configurations, we present the latency for the Swin-MoE-Base model under different batch sizes in Figure 7.
+When the workload is relatively small, a model-centric setting is more efficient, while a data-centric setting becomes more efficient with a relatively large workload.
+Experiments are all conducted on the homogeneous machine with automatic mixed precision in PyTorch.
+Figure 7: Latency comparison with data-centric and model-centric configurations.
+We record the average latency per step with 2k steps in total.
+The data-centric setting presents a more gentle trend when scaling workload.
+Heterogeneous Experiments.
+We demonstrate the heterogeneous awareness of our HEXA-MoE via experiments on different devices under both data-centric and model-centric configurations, and visualize the average latency with different division strategies in Figure 8.
+Since the $2$ GPUs in our heterogeneous machine have similar computing capacities, we also adjust the power limit for each device.
+We first examine the computing capacity for these experimental settings, and provide the results in Table 2.
+The details of the proxy task are provided in Appendix C.
+In Figure 8, we adjust the batch size on each device under data-centric configuration.
+Essentially the optimal proportion is very close to the examined computing capacity proportion, and it has a certain deviation from uniform division when the power constraints of the two devices are not equal, while the optimal division is also uniform division when the power constraints are equal.
+Specifically, when setting the power constraint for $D_{0}$ to be 100W and $D_{1}$ to be 300W, employing our heterogeneous-aware allocation can make the average latency to be 13.2% lower than naive division, and when setting $D_{0}$ to be 300W and $D_{1}$ to be 100W, our approach can reduce the average latency by 25.3%.
+In Figure 8, we adjust the allocated sub-dimension proportion of the hidden size for each MoE layer on each device.
+Adapting the sub-dimension proportion to be close to the computing capacity proportion of heterogeneous devices can also substantially minimize average latency, similar to the data-centric setting.
+Specifically, when setting the power constraint of $D_{0}$ to be 100W and $D_{1}$ to be 300W, our approach can achieve a 6.3% reduction on average latency compared to uniform division, and when setting $D_{0}$ to be 300W and $D_{1}$ to be 100W, we can reduce it by 11.9%.
+Although the reduction is not as significant as a data-centric setting, we can still achieve a relative speed-up.
+These demonstrate that our HEXA-MoE can essentially maximize the utilization of heterogeneous computing resources.
+Table 2: Examining computing capacity proportion for heterogeneous devices. $P$, $T$ and $R$ denote power constraint (W), average latency (s), and computing capacity proportion.
+Case 1 Case 2 Case 3
+$P$ $T$ $R$ $P$ $T$ $R$ $P$ $T$ $R$
+$D_{0}$ 100 4.58 0.40 300 3.20 0.50 300 3.28 0.74
+$D_{1}$ 300 3.06 0.60 300 3.18 0.50 100 9.42 0.26
+Figure 8: Latency analysis on heterogeneous devices with both data-centric and model-centric configuration.
+Experiments are all conducted with Swin-MoE-Small model using 4 global experts and top-2 routing.
+The average latency falls minimal when the division proportion is close to the computing capacity proportion in each case.
+4.3 Ablation Studies
+We examine the effectiveness of each component for HEXA-MoE via measuring the impact of expert-specific operators, pipeline-shared cache, fused kernel, data- & model-centric and memory optimization, respectively.
+$2$ metrics are evaluated, including average latency and memory footprint.
+We take distributed training of Swin-MoE-Base model on homogeneous devices as benchmark for ablation studies.
+Since dispatch & combine operations are inevitable if we dispense with our expert-specific operators, we directly take the performance of Tutel in this case.
+Memory Footprint Breakdown.
+We take 8 experts, top-4 routing, and batch size 40 as an example to break down memory footprint, and visualize the results in Figure 9. We can draw some insights from it:
+  1. $\star$
+    The employment of pipeline-shared cache slightly increases memory footprint, while expert-specific fused kernel has no impact.
+  2. $\star$
+    Pipeline-shared cache can effectively reduce memory footprint under a data-centric setting, since the memory footprint would surpass Tutel without it.
+  3. $\star$
+    HEXA-MoE can reduce memory footprint compared to baseline even without memory optimization, and optimizing it can further reduce memory footprint.
+Latency Breakdown.
+We set 4 experts, top-4 routing, and batch size 80 to break down training latency, and visualize results in Figure 9. We can draw some insights as:
+  1. $\star$
+    The employment of expert-specific fused kernel, data-centric setting, and communication-computation overlap can effectively reduce latency.
+Meanwhile, the combination of data-centric setting and communication-computation overlap can further empower our HEXA-MoE with notably reduced latency.
+  2. $\star$
+    Expert-specific operators play a significant part in speeding up MoE computing since the latency for Tutel is much longer than our model-centric setting.
+  3. $\star$
+    Although pipeline-shared cache and memory optimization can slightly increase latency, they can also lead to significant reduction on memory footprint, as demonstrated in memory footprint breakdown.
+Figure 9: Memory and latency breakdown.
+We analyze the effectiveness of our proposed pipeline-shared cache, fused kernel, and memory optimization.
+We take Tutel as a baseline and visualize the differences with it.
+5 Conclusion
+We introduce HEXA-MoE, a completely static MoE training framework with minimal memory footprint and heterogeneous-awareness, materialized by the proposed expert-specific operators and tensor parallelism.
+We provide specialized GPU kernels, which can both save memory footprint and reduce overall latency.
+For different workload scales, we provide data- and model-centric configurations for enhanced efficiency, and propose a memory-efficient communication-computation overlapping scheme to overcome the shortcomings of previous work.
+Homogeneous experiments show that HEXA-MoE can save $10\%$-$48\%$ memory footprint while achieving $0.5$-$4.3$$\times$ speed up compared to state-of-the-art MoE libraries.
+Heterogeneous experiments show that HEXA-MoE can substantially minimize runtime and better utilize global computing resources by employing optimal parallel configuration.
+HEXA-MoE endows MoE computing with an in-place manner, where routing choice would yield almost no impact on hardware workload, which may inspire further research on algorithm designs of MoE.
+Impact Statement
+This paper aims to improve the computation efficiency for training MoE models.
+The efficiency advantage of such models might help democratize access of large-scale foundation models.
+On the other hand, whether such new algorithm would affect known issues such as biased and harmful outputs of language models remains an unexplored research question.
+Appendix A Principles of CUDA Programming
+Introduction to CUDA Programming:
+A GPU program can be viewed as a thread grid in Figure 10(a), which is $3$-dimensional and contains massive thread blocks.
+A thread block is also $3$-dimensional and contains up to $1024$ threads.
+Computation is mapped to these threads and implemented in parallel.
+Physically a GPU is composed of massive streaming multiprocessors (SM), where we show the memory hierarchy in Figure 10(b).
+Each SM loads data from global memory to its registers for parallel computing, and caching with shared memory can be faster.
+An SM contains many processing units including CUDA core and Tensor core.
+The former is used for general parallel computing, while the latter is specialized for matrix multiplication with mixed precision, which is first introduced in NVIDIA Volta architecture (Choquette et al., 2018).
+During execution, an SM is given one or more thread blocks, which are partitioned into warps and each warp gets scheduled by a warp scheduler.
+Mathematical Symbols:
+We provide the description of mathematical symbols used in this paper in Table 3, including variables in MoE computing and constant values in CUDA programming.
+To utilize Tensor core for faster matrix processing, we have to call the nvcuda::wmma (warped matrix multiplication and add) interface to compute a fixed-sized matrix multiplication such as $\times 16$ in a single warp, therefore the number of threads in a block has to be divisible by the constant value WARP.
+In the CUDA implementation of our method, a thread block loads BLK tokens each time from global memory to shared memory to conduct the parallelized computation.
 Table 3: Description of symbols.
@@
 $D_{i},D_{o}$ Input and output size of the FFN.
-$\bm{x}$ Data batch with $N$ tokens.
-$\bm{x}^{e},N_{e}$ Tokens in $\bm{x}$ routed to expert $e$ with number $N_{e}$.
-$\{\mathcal{R}_{i}(\bm{x})\}_{i=0}^{k-1}$ Routing choice for $\bm{x}$ under top-$k$ routing.
-$\bm{W_{1}},\bm{W_{2}}$ Weights for 2 MLPs, shaped as $(E,D_{1},D_{2})$.
-$\bm{b_{1}},\bm{b_{2}}$ Biases of the two MLPs, shaped as $(E,D)$.
+$\boldsymbol{x}$ Data batch with $N$ tokens.
+$\boldsymbol{x}^{e},N_{e}$ Tokens in $\boldsymbol{x}$ routed to expert $e$ with number $N_{e}$.
+$\{\mathcal{R}_{i}(\boldsymbol{x})\}_{i=0}^{k-1}$ Routing choice for $\boldsymbol{x}$ under top-$k$ routing.
+$\boldsymbol{W_{1}},\boldsymbol{W_{2}}$ Weights for 2 MLPs, shaped as $(E,D_{1},D_{2})$.
+$\boldsymbol{b_{1}},\boldsymbol{b_{2}}$ Biases of the two MLPs, shaped as $(E,D)$.
 BLK Block size for expert-specific operators.
@@
 TIMES A thread block has WARP×$\times$×TIMES threads.
-2 Related Works
-Scaling Transformers with MoE.
-Transformer models can be scaled up via the Mixture-of-Experts (MoE) for better learning capacity and generalization while enjoying a sub-linear increase in computation overhead due to its sparsity.
-It has been proven in many research fields such as natural language processing Fedus et al.
-(2022); Jiang et al.
-(2024); Du et al.
-(2022), machine vision Riquelme et al.
-(2021); Fan et al.
-(2022) and multi-modal learning Bao et al.
-(2022).
-Open-Source MoE Computing Libraries.
-FastMoE He et al.
-(2021) pioneered the first PyTorch open-source implementation for distributed MoE.
-Based on it, FasterMoE He et al.
-(2022a) proposes dynamic shadowing and smart scheduling to tackle load imbalance and improve parallelism.
-Tutel Hwang et al.
-(2023) implements switchable parallelism and dynamic pipelining, improving adaptability and scalability.
-MegaBlocks Gale et al.
-(2023) proposes block-sparse operations and corresponding GPU kernels to mitigate dynamic routing issues in MoEs.
-Improving efficiency for MoE computing.
-MoE computing faces significant challenges in communication and memory.
-Recent research has explored alleviating them from various perspectives: Lina Li et al.
-(2023) proposes to dynamically schedule resources to reduce latency and tackle the all-to-all bottleneck.
-Janus Liu et al.
-(2023) proposes a data-centric paradigm that replaces traditional all-to-all communication with asynchronous expert fetching, enabling overlapped communication and computation for enhanced parallelism.
-ScheMoE Shi et al.
-(2024) proposes a generic scheduling framework for optimal communication and computation scheduling during MoE training.
-MPMoE Zhang et al.
-(2024) accelerates MoE training through adaptive and memory-efficient pipeline parallelism.
-SmartMoE Zhai et al.
-(2023) proposes an efficient searching algorithm to identify optimization opportunities within an expanded hybrid parallelism space, tailored for data-sensitive MoE models.
-PipeMoE Shi et al.
-(2023) adaptively pipelines communications and computations in MoE to mask communication latency.
-It also provides an optimal strategy to determine pipeline degree to minimize overall iteration time.
-3 Preliminary
+(a) Thread hierarchy for CUDA programming.
+(b) Memory hierarchy for a GPU.
 Figure 10: Thread hierarchy for CUDA programming and memory hierarchy for a typical NVIDIA GPU.
-Introduction to CUDA Programming:
-A GPU program can be viewed as a thread grid in Figure LABEL:fig:thread-hierarchy, which is $3$-dimensional and contains massive thread blocks.
-A thread block is also $3$-dimensional and contains up to $1024$ threads.
-Computation is mapped to these threads and implemented in parallel.
-Physically a GPU is composed of massive streaming multiprocessors (SM), where we show the memory hierarchy in Figure LABEL:fig:memory-hierarchy.
-Each SM loads data from global memory to its registers for parallel computing, and caching with shared memory can be faster.
-An SM contains many processing units including CUDA core and Tensor core.
-The former is used for general parallel computing, while the latter is specialized for matrix multiplication with mixed precision, which is first introduced in NVIDIA Volta architecture Choquette et al.
-(2018).
-During execution, an SM is given one or more thread blocks, which are partitioned into warps and each warp gets scheduled by a warp scheduler.
-Mathematical Symbols:
-We provide the description of mathematical symbols used in this paper in Table 1, including variables in MoE computing and constant values in CUDA programming.
-To utilize Tensor core for faster matrix processing, we have to call the nvcuda::wmma (warped matrix multiplication and add) interface to compute a fixed-sized matrix multiplication such as $\times 16$ in a single warp, therefore the number of threads in a block has to be divisible by the constant value WARP.
-In the CUDA implementation of our method, a thread block loads BLK tokens each time from global memory to shared memory to conduct the parallelized computation.
-Figure 5: Comparison between conventional and expert-specific formulation for MoE computing.
-We take top-1 routing for illustration and present the corresponding relation of each formula in the MoE forward and backward propagation.
-4 Method
-We start our exploration by first formulating the forward and backward propagation of a single MoE layer with and without GeMM, which induces the same calculation results in different ways (Sec.
-4.1).
-Specifically, GeMM-based MoE computing depends on inefficient dispatch & combine operations, while our expert-specific operators serve as an alternative, performing in-place and inducing ZERO computation redundancy.
-After that, we expound the details for our optimized CUDA implementation (Sec.
-4.2).
-We also introduce a novel expert-specific fused operator in this part, specialized for parallel MoE backward propagation.
-These empower our HEXA-MoE with high computation efficiency.
-Next, we consider different workload scales, and adapt HEXA-MoE efficiently to both data- and model-centric settings (Sec.
-4.3).
-Finally, we adapt HEXA-MoE to heterogeneous devices, and provide an expert allocation approach to minimize the average latency so as to utilize better heterogeneous computing capacities (Sec.
-4.4).
-4.1 MoE Computing Formulation
-GeMM-based MoE Computing.
-Current MoE libraries are mainly built on expert parallelism Fedus et al.
-(2022) and dispatch & combine operations Hwang et al.
-(2023).
-Specifically, “expert parallelism” means distributing the experts to different devices, “dispatch” means assigning the experts with different tokens based on the routing choice, and “combine” means restoring the original data batch from dispatched tokens.
-In conventional settings, all-to-all communication is required to further assign the dispatched tokens to the corresponding device, where the corresponding expert is kept.
-To employ GeMM on each device, we need to calibrate the workload of each expert via token padding or discarding and build a new data batch.
-Taking top-1 routing as an example, we formulate the forward and backward propagation in Figure 5, denoting $\ell$ as loss value and $\bm{x}^{e}$ as tokens dispatched to expert $e$ (with $N_{\textrm{e}}$ tokens).
-In backward propagation, the gradients of the output have been provided by the auto-differentiation program.
-Based on dispatch and combine operations, all variables can be derived via basic matrix operations such as summation and multiplication.
-Although we can utilize the high-performance GeMM interface, these operations are inefficient, since the workload of each expert varies dynamically in each step, and token padding or discarding has to be employed to construct new data batches for local experts.
-It can be overcome by our expert-specific design.
-MoE Computing via Expert-Specific Operators.
-Based on the above formulations, we propose to re-implement in an in-place manner to address the inefficient teaser of dispatch & combine operations.
-Specifically, we find that the forward and backward propagation of a single MoE layer can be reformulated as built upon 3 specialized operators from an expert-specific perspective, namely expert-specific operators.
-We take top-$1$ routing for illustration, while formulations for top-$k$ routing are provided in Appendix A.
-Figure 9: Illustration of the proposed operators.
-Take $10$ tokens, $4$ global experts, and block size $4$ as an example.
-For ESTMM, the $2$ input batches are in a re-indexed format, while for others both the raw batch and re-index vector are provided.
-  1.
-Expert-specific matrix multiplication (ESMM): given input batch $\bm{x}$ with $N$ tokens, routing choice $\mathcal{R}(\bm{x})$, weight $\bm{W}$ and bias $\bm{b}$, we formulate it as $\bm{y}=\textit{ESMM}(\bm{x},\bm{W},\bm{b},\mathcal{R}(\bm{x}))$, where $\bm{y}_{i}$ is derived from $\bm{x}_{i}$, as well as the corresponding $\bm{W}_{\mathcal{R}(\bm{x}_{i})}$ and $\bm{b}_{\mathcal{R}(\bm{x}_{i})}$.
-ESMM is applied to get the hidden and output tokens in forward, and the gradients of hidden and input tokens in backward propagation, as shown in Figure LABEL:fig:esmm.
-  2.
-Expert-specific summation (ESS): given input batch $\bm{x}$ with $N$ tokens and routing choice $\mathcal{R}(\bm{x})$, it is formulated as $\bm{y}=\textit{ESS}(\bm{x},\mathcal{R}(\bm{x}))$, where tokens routed to expert $e$ are accumulated and recorded in $\bm{y}[e]$.
-ESS is applied to get the gradients of bias in backward propagation, as illustrated in Figure LABEL:fig:ess.
-  3.
-Expert-specific transposed matrix multiplication (ESTMM): the inputs include two batches $\bm{x}_{1}$ and $\bm{x}_{2}$ (both with $N$ tokens), and the routing choice $\mathcal{R}(\bm{x})$.
-Both $\bm{x}_{1}$ and $\bm{x}_{2}$ are the ESMM result of $\bm{x}$, thus sharing the same routing choice.
-It is formulated as $\bm{y}=\textit{ESTMM}(\bm{x}_{1},\bm{x}_{2},\mathcal{R}(\bm{x}))$.
-For the $i$-th channel of $\bm{x}_{1}$ and $j$-th channel of $\bm{x}_{2}$, we first prepare a zero vector $\bm{c}$ with $E$ elements, and accumulate $\bm{x}_{1}[m,i]\cdot\bm{x}_{2}[m,j]$ to $\bm{c}[\mathcal{R}(\bm{x}_{m})]$ for all $0\leq m<N$.
-After that we assign $\bm{c}[k]$ to $\bm{y}[k,i,j]$ for all $0\leq k<E$.
-ESTMM is applied to get the gradients of weights in backward propagation, as illustrated in Figure LABEL:fig:estmm.
-Based on these expert-specific operators, we can now implement MoE computing in a novel in-place manner, as formulated in Figure 5.
-We compare each stage between our method and conventional formulation, and the output of each token is ensured to be consistent for both cases.
-In forward propagation, ESMM performs as an alternative to common matrix multiplication, while in backward propagation, the gradient of $\bm{y}$ is also provided by the auto-differentiation program.
-ESS performs as an alternative to tensor summation, while ESMM and ESTMM are designed as $2$ specialized cases for matrix multiplication from the expert-specific perspective.
-With the help of our expert-specific design, we can now directly compute the output for all MoE computing tasks rather than dispatching the tokens to each expert before MoE computing, introducing zero redundancy.
-4.2 Optimized CUDA Implementations
-Figure 13: Naive expert-specific MoE computing and the memory-efficient version.
-We take top-2 routing and 4 tokens as an example.
-In the naive version, we have to compute from top-$1$ to top-$k$ respectively, write each result back, and sum them up for the final output, while a memory-efficient version adopts AtomicAdd instead, saving much memory.
-We also present a comparison for the 2 schemes, assuming the memory for one input & output token as 1 and hidden token as 4.
-Re-Index based Expert-Specific Operators.
-Although expert-specific operators enable MoE computing to be performed in an in-place manner, we should not implement them directly, since simply computing for each token in parallel cannot fully utilize the locality of GPU, which restricts the performance.
-Considering that we cannot guarantee the adjacent tokens to be routed to the same expert, we have to introduce other tensors as auxiliarys.
-In HEXA-MoE, we additionally adopt a re-index vector, as illustrated in Figure LABEL:fig:re-index.
-Specifically, we re-organize the token indices based on the routing choice, i.e., gathering indices for tokens that routed to the same expert into a sub-vector, and padding it with -1 to make it divisible by a constant value BLK.
-The algorithm details are provided in Appendix A.
-We present the details of optimized ESMM implementation in Figure LABEL:fig:esmm, where a thread block loads BLK input tokens each time based on the re-index vector, as well as the corresponding expert parameters.
-Since these BLK tokens are routed to the same expert, we only need to load one expert each time.
-We then accumulate the dot product results along the dimension of the input feature and write the result for each token back to its position in global memory, also based on the re-index vector.
-The details of optimized ESS is shown in Figure LABEL:fig:ess, where each thread block is employed to compute some channels of one expert, with BLK tokens one time.
-After accumulating all tokens within one expert, it writes the summation back to global memory.
-ESTMM is illustrated in Figure LABEL:fig:estmm, where we present the input batches in a re-indexed manner.
-Specifically, the two input batches are the ESMM results of the same tensor, sharing the same re-index vector.
-Each thread block loads BLK channels of both batches for tokens that are routed to the same expert.
-After accumulation, results are written back to the corresponding positions respectively.
-We also provide algorithm details of this part with PyTorch style pseudo-code in Appendix A.
-Memory Optimization for Top-$k$ Routing.
-We find that when extending the routing choice from top-$1$ to top-$k$, the memory consumption would increase significantly.
-To tackle this issue, we present a simplified pipeline in Figure 13, where only input, hidden, and output tokens in one MoE layer are provided.
-For naive forward propagation in Figure LABEL:fig:forward-normal and backward in Figure LABEL:fig:backward-normal, we have to compute the result of all the $k$ routing choices and sum them up to obtain the final output.
-Specifically, in forward propagation, we have to sum up the outputs of ESMM for the second MLP layer to get the output tokens, while in backward propagation we have to sum up the outputs of ESMM for the first MLP layer to get the gradients for input tokens.
-This leads to extra memory allocation for pre-summed tensors, including pre-summed output tokens in forward and pre-summed input tokens in backward.
-To optimize memory consumption when scaling from top-1 routing to top-$k$, we only allocate memory for the hidden tokens, and sum up the result of each routing choice via atomicAdd in CUDA programming, as illustrated in Figure LABEL:fig:memory-efficienct.
-This can result in obvious memory reduction, as we presented in Figure LABEL:fig:mem-opt.
-Figure 15: Visualization of the training pipeline and shared cache under the data-centric setting.
-Before all gather communication, each device loads the kept parameter chunk to its shared cache, after that each cache will be filled with the whole parameters of an MoE layer.
-Figure 16: Tensor parallelism for HEXA-MoE under model-centric setting.
-Take $2$ GPUs, $1$ MoE layer, and $4$ experts as an example, assuming the local batch contains $4$ tokens.
-We refactor classical tensor parallelism with expert-specific operators.
-Fused Kernel in Backward Propagation.
-For a single multi-layer perception (MLP) $\bm{y}=\bm{x}\cdot\bm{W}+\bm{b}$, the gradients for $\bm{x}$, $\bm{W}$ and $\bm{b}$ can be computed in parallel, as they are unrelated with each other.
-Therefore, we can integrate the implementation of ESS, ESTMM, and ESMM together into one kernel function named expert-specific fused kernel (ESFK) for better parallelism.
-However, directly integrating them is hard to implement, since the shape of thread blocks required for each operator and the implementation details for each operator vary a lot with each other.
-To this end, we set the shape of the thread block to be the same for each operator, and expand one dimension for ESMM and ESS so that the thread grids are all $3$-dimensional.
-Taking a single FFN layer under top-$1$ routing as an example, where we present the shape of allocated thread block and thread grid for both forward and backward propagation in Table 6 in Appendix B.
-By transposing the shape or expanding one dimension for the thread grid of an operator, we can align the 1st and 2nd dimensions while summing the 3rd dimension together for an integrated grid.
-In this way, the backward propagation for one FFN can be re-implemented with only $2$ fused kernel function and $1$ element-wise dot production.
-4.3 Parallel Strategy for Different Workload Scale
-Data Parallelism for Data-Centric Setting.
-Data-centric is a practical and efficient approach for both training and inference when the scale of the data batch is larger than the model parameters.
-HEXA-MoE distributes MoE parameters to different devices along the hidden size dimension of FFN, based on tensor parallelism.
-In a data-centric setting, each device gathers the whole MoE parameters from other devices and performs computing for each MoE layer locally, similar to data parallelism.
-Although MoE computing under data-centric configuration has been explored by Janus Liu et al.
-(2023), memory efficiency has not yet been fully studied.
-Janus pre-fetchs the required MoE parameters for each layer progressively in forward propagation and preserves all of them locally for backward propagation so that no communication is required when computing for gradients.
-However, the sheer size of MoE parameters can lead to a significant increase in memory usage, making it inefficient and impracticable for wider deployment.
-To tackle this issue, we propose to allocate an additional region on the GPU memory of each device, which is adopted to dynamically cache the gathered MoE parameters for each pipeline stage, named as pipeline-shared cache.
-Since each device keeps a subset of the FFN hidden dimension for all experts in each layer, we need to employ all gather communication among devices before the computing of each MoE layer, preparing for the required parameters during both forward and backward propagation, as illustrated in Figure LABEL:fig:pipeline and LABEL:fig:cache.
-To achieve better parallelism, all gather communication can be overlapped with other operations such as attention and router, as shown in Figure LABEL:fig:pipeline.
-In this way, each device would not have to preserve all the MoE parameters for backward propagation, while communication overhead can also be well masked, therefore both memory efficiency and computing efficiency can be achieved.
-Tensor Parallelism for Model-Centric Setting.
-In the case that the scale of data is less than model parameters, model-centric configuration turns out to be more efficient than data-centric.
-To distribute MoE parameters on multiple devices, we modify the classical tensor parallelism with our proposed expert-specific operators, as shown in Figure 16.
-All gather communication is employed to synchronize local data batches before and after each MoE layer, and a pipeline-shared cache is not required in this setting.
-In Figure 2 we present a comparison between expert parallelism and tensor parallelism for the MoE layer, where HEXA-MoE takes small modification based on standard tensor parallelism Narayanan et al.
-(2021).
-We distribute each MoE layer among different devices along the hidden size dimension for each expert, and during MoE computing, each device computes the all gathered data batches with only the local MoE parameter chunk using ESMM, and the output tokens are all reduced with sum operation in forward propagation.
-During backward propagation, all gather and all reduce communications are interchanged, while ESMM are replaced by ESMM, ESS, and ESTMM to derive the gradients for input tokens, bias, and weights, respectively, which can also be replaced by the fused operator ESFK.
-Figure 18: Average memory usage for training Swin-Transformer-MoE models.
-We take 8 global experts and examine all cases from top-1 to top-8 routings.
-Experiments are conducted on 2 homogeneous GPUs using automatic mixed precision in PyTorch.
-The batch size is set to 40 for all cases.
-We record the average GPU memory consumption (GB) on each device.
-4.4 Heterogeneous-aware Expert Allocation
-Workload Division for a different configuration.
-For data-centric configuration, since it can be essentially viewed as data parallelism, the workload mainly depends on batch size.
-For model-centric configuration, the workload of each device for the MoE layer mainly depends on the allocated sub-dimension of the hidden size with each expert, as we have replaced conventional expert parallelism with pure tensor parallelism with the aid of expert-specific operators.
-Practical Expert Allocation on Heterogeneous Devices.
-To adapt our HEXA-MoE to heterogeneous devices, we propose a specialized expert allocation approach that better utilizes the overall computing capacity by adjusting the workload on each device.
-Specifically, we have to first examine the computing capacity of each device by averaging its latency on the same proxy task of heavy computing, such as large-scale matrix multiplication or large-scale MLP computing.
-Denoting the results as $\{t_{i}\}_{i=0}^{N-1}$ for $N$ devices that distribute the MoE model.
-For the data-centric setting, we assign different local batch size $\{B_{i}\}_{i=0}^{N-1}$ to different devices based on the observed latency, as shown in Equation 1:
-$B_{i}=\frac{1/t_{i}}{\sum_{j=0}^{N-1}1/t_{j}}\cdot B_{\textrm{global}},$ (1)
-where we denote $B_{\textrm{global}}$ as the global batch size, i.e., $B_{\textrm{global}}=\sum_{j=0}^{N-1}B_{i}$.
-For the model-centric setting, we denote the sub-dimension of hidden size for one MoE layer on each device as $\{h_{i}\}_{i=0}^{N-1}$ with total hidden size $H$.
-The assigned sub-dimension $h_{i}$ on device $i$ can be derived as Equation 2.
-$h_{i}=\frac{1/t_{i}}{\sum_{j=0}^{N-1}1/t_{j}}\cdot H.$ (2)
-To ensure that both $\{B_{i}\}_{i=0}^{N-1}$ and $\{h_{i}\}_{i=0}^{N-1}$ are integers, we need to further round them up or down while making sure that the summation is exactly $B_{\textrm{global}}$ or $H$.
-5 Experiments
-5.1 Experimental Setup
-We implement all the experiments using 2 machines, namely $M_{\textrm{homo}}$ and $M_{\textrm{hete}}$.
-Specifically, $M_{\textrm{homo}}$ is composed of 4 homogeneous GPUs, while $M_{\textrm{hete}}$ contains 2 heterogeneous GPUs.
-Details for the machines and GPUs are provided in Table 2.
-We take the training process of Swin-Transformer-MoE as the benchmark for all the experiments, including memory analysis, latency analysis, heterogeneous computing analysis, and ablation studies.
-Apart from heterogeneous computing analysis, all the experiments are conducted on $M_{\textrm{homo}}$.
-We also adopt both the Small and Base scales for the Swin-MoE model, following Tutel Hwang et al.
-(2023).
-Experiments are conducted with PyTorch, taking nccl as the communication backend.
-All the experiments are implemented in a distributed manner based on the distributed data-parallel interface in PyTorch, and batch size is the batch size on a single device rather than the global workload.
-Figure 19: Average latency for training Swin-Transformer-MoE models.
-Experiments are conducted on 4 homogeneous GPUs with 4 global experts.
-We set different batch sizes for different models under different routing strategies to maximize the utilization of GPU memory.
-We record the average latency for one step (s) during training with 2k steps in total.
-5.2 Experimental Results
-Memory Analysis on Homogeneous Devices.
-We analyze the memory consumption for different MoE libraries on homogeneous devices.
-We adopt 8 global experts and examine all cases from top-$1$ routing to top-$8$.
-Results are visualized in Figure 18.
-HEXA-MoE can reduce $10\%$-$48\%$ memory usage compared to Tutel Hwang et al.
-(2023) and MegaBlocks Gale et al.
-(2023), and the memory usage for model-centric setting is slightly less than data-centric setting due to the employment of pipeline-shared cache.
-When scaling top-$1$ routing to top-$k$ routing, the memory reduction becomes more significant, since we only enlarge the memory allocation for the hidden tokens, as illustrated in Figure 13.
-Meanwhile, the trend of memory consumption increase for our method is also much more gentle than Tutel and MegaBlocks when scaling from top-1 routing to top-$k$, owing to the introduction of memory-latency trade-off.
-Since HEXA-MoE performs MoE computing in an in-place manner and introduces no token padding, we can still save a certain amount of memory under top-$1$ routing.
-Table 2: Details of the machines and GPUs used for both homogeneous and heterogeneous experiments.
-Notation Number & Hardware Specs Memory
-CPU $M_{\textrm{homo}}$ 2×2$\times$2 × Intel Xeon Platinum 8352V 2.10GHz 1008 GB
-$M_{\textrm{hete}}$ 2×2$\times$2 × Intel Xeon Gold 6130 2.10GHz 62.5 GB
-GPU $D$ (in $M_{\textrm{homo}}$) 4×4$\times$4 × NVIDIA GeForce RTX 4090 24 GB
-$D_{0}$ (in $M_{\textrm{hete}}$) 1×1$\times$1 × NVIDIA TITAN RTX 24 GB
-$D_{1}$ (in $M_{\textrm{hete}}$) 1×1$\times$1 × NVIDIA GeForce RTX 2080 Ti 11 GB
-Latency Analysis on Homogeneous Devices.
-We analyze the average latency for one training step, and present the results in Figure 19.
-To fully utilize the GPU memory, we set different batch sizes for different routing configurations and different model scales.
-HEXA-MoE achieves $0.5$-$4.3$$\times$ speed up compared to Tutel Hwang et al.
-(2023) and MegaBlocks Gale et al.
-(2023).
-Since the scale of workload is larger than MoE parameters in our experiments, the average latency under a data-centric setting is less than model-centric in these cases, and the advantage of HEXA-MoE turns more obvious while enlarging batch size.
-To better visualize the latency comparison between data-centric and model-centric configurations, we present the latency for the Swin-MoE-Base model under different batch sizes in Figure 20.
-When the workload is relatively small, a model-centric setting is more efficient, while a data-centric setting becomes more efficient with a relatively large workload.
-Experiments are all conducted on the homogeneous machine with automatic mixed precision in PyTorch.
-Figure 20: Latency comparison with data-centric and model-centric configurations.
-We record average latency per step with 2k steps in total.
-The data-centric setting presents a more gentle trend when scaling workload.
-Heterogeneous Experiments.
-We demonstrate the heterogeneous awareness of our HEXA-MoE via experiments on different devices under both data-centric and model-centric configurations, and visualize the average latency with different division strategies in Figure 22.
-Since the $2$ GPUs in our heterogeneous machine have similar computing capacities, we also adjust the power limit for each device.
-We first examine the computing capacity for these experimental settings, and provide the results in Table 3.
-The details of the proxy task are provided in Appendix B.
-In Figure LABEL:subfig:data-centric-hete, we adjust the batch size on each device under data-centric configuration.
-Essentially the optimal proportion is very close to the examined computing capacity proportion, and it has a certain deviation from uniform division when the power constraints of the two devices are not equal, while the optimal division is also uniform division when the power constraints are equal.
-Specifically, when setting the power constraint for $D_{0}$ to be 100W and $D_{1}$ to be 300W, employing our heterogeneous-aware allocation can make the average latency to be 13.2% lower than naive division, and when setting $D_{0}$ to be 300W and $D_{1}$ to be 100W, our approach can reduce the average latency by 25.3%.
-In Figure LABEL:subfig:model-centric-hete, we adjust the allocated sub-dimension proportion of the hidden size for each MoE layer on each device.
-Adapting the sub-dimension proportion to be close to the computing capacity proportion of heterogeneous devices can also substantially minimize average latency, similar to the data-centric setting.
-Specifically, when setting the power constraint of $D_{0}$ to be 100W and $D_{1}$ to be 300W, our approach can achieve a 6.3% reduction on average latency compared to uniform division, and when setting $D_{0}$ to be 300W and $D_{1}$ to be 100W, we can reduce it by 11.9%.
-Although the reduction is not as significant as a data-centric setting, we can still achieve a relative speed-up.
-These demonstrate that our HEXA-MoE can essentially maximize the utilization of heterogeneous computing resources.
-Table 3: Examining computing capacity for heterogeneous devices. $P$, $T$ and $R$ denote power constraint (W), average latency (s), and computing capacity proportion.
-Case 1 Case 2 Case 3
-$P$ $T$ $R$ $P$ $T$ $R$ $P$ $T$ $R$
-$D_{0}$ 100 4.58 0.40 300 3.20 0.50 300 3.28 0.74
-$D_{1}$ 300 3.06 0.60 300 3.18 0.50 100 9.42 0.26
-Figure 22: Latency analysis on heterogeneous devices with both data-centric and model-centric configuration.
-Experiments are all conducted with Swin-MoE-Small model using 4 global experts and top-2 routing.
-The average latency falls minimal when the division proportion is close to the computing capacity proportion in each case.
-5.3 Ablation Studies
-We examine the effectiveness of each component for our HEXA-MoE via measuring the impact of expert-specific operators, pipeline-shared cache, fused kernel, data- & model-centric and memory-latency trade-off respectively on two indicators, i.e., average latency and memory usage.
-We take the distributed training of the Swin-MoE-Base model on homogeneous devices as the benchmark for ablation study experiments.
-Since dispatch & combine operations are inevitable if we dispense with our expert-specific operators, we directly take the performance of Tutel in this case.
-Memory Usage Breakdown.
-We take 8 experts, top-4 routing, and batch size 40 as an example to break down memory usage, and visualize the results in Figure LABEL:fig:mem_breakdown. We can draw some insights from it:
-  1. $\star$
-    The employment of pipeline-shared cache slightly increases memory consumption, while expert-specific fused kernel has no impact.
-  2. $\star$
-    Pipeline-shared cache can effectively reduce memory consumption under a data-centric setting since the memory usage would surpass Tutel without it.
-  3. $\star$
-    Our HEXA-MoE can reduce memory usage compared to baseline even without memory optimization, and optimizing it can further reduce memory consumption.
-Latency Breakdown.
-We set 4 experts, top-4 routing, and batch size 80 to break down training latency, and visualize results in Figure LABEL:fig:latency_breakdown. We can draw some insights as:
-  1. $\star$
-    The employment of expert-specific fused kernel, data-centric setting, and communication-computation overlap can effectively reduce latency.
-Meanwhile, the combination of data-centric setting and communication-computation overlap can further empower our HEXA-MoE with notably reduced latency.
-  2. $\star$
-    Expert-specific operators play a significant part in speeding up MoE computing since the latency for Tutel is much longer than our model-centric setting.
-  3. $\star$
-    Although the employment of pipeline-shared cache and memory optimization would slightly increase latency, they can also lead to significant memory reduction, as demonstrated in memory usage breakdown.
-Figure 24: Memory and latency breakdown.
-We analyze the effectiveness of our proposed pipeline-shared cache, fused kernel, and memory-latency trade-off.
-We take Tutel as a baseline and visualize the differences with it.
-6 Conclusion
-In this paper, we analyze the limitations of current MoE libraries in detail and propose novel expert-specific operators as an alternative to GeMM so that MoE computing can be performed in an in-place manner with zero computation redundancy.
-We reformulate the forward and backward propagation of the MoE layer with these operators and replace conventional expert parallelism with tensor parallelism, i.e., splitting the parameters of the MoE layer along the hidden size dimension to distribute them on different devices.
-We provide GPU implementations for these operators in a highly parallelized manner, which can both save memory consumption and reduce overall latency.
-For different workload scales, we implement the MoE model under both data- and model-centric configurations.
-To tackle the teaser of heavy memory consumption in previous data-centric works, we propose a pipeline-shared cache for better memory efficiency.
-Homogeneous experiments show that HEXA-MoE can save $10\%$-$48\%$ memory usage while achieving $0.5$-$4.3$$\times$ speed up compared to state-of-the-art MoE libraries.
-To adapt HEXA-MoE to heterogeneous devices, we propose to assign different batch sizes or sub-dimensions of hidden sizes for different devices based on the heterogeneous computing capacity.
-Experiments show that our adaption can substantially minimize average latency and better utilize the whole computing capacity.
-Meanwhile, our HEXA-MoE empowers MoE computing to be conducted in an in-place manner, i.e., the routing choice would yield almost no impact on hardware workload, which can facilitate further research on MoE algorithms.
-Appendix A Algorithm Details
-We present the algorithm details for re-index vector construction as well as the expert-specific operators, and taking top-1 routing as an example for illustration, shown in Algorithm 1, 3, 2 and 4.
-In Algorithm 1, we re-arrange the routing choice vector $\mathcal{R}$ into re-indexed token vector $v$, along with the token index starting vector $idx$, which satisfies $idx[0]=0$ and $idx[E]=N^{\prime}$.
-We provide the length and range for the vectors in Table 4.
-Table 4: Explanation for the auxiliary vectors.
-vector length range
-$\mathcal{R}$ $N$ $0\leq\mathcal{R}[i]<E$ for $0\leq i<N$
-$v$ $N^{\prime}$ $0\leq v[i]<N$ for $0\leq i<N^{\prime}$
-$idx$ $1+E$ $0\leq idx[i]<N^{\prime}$ for $0\leq i<1+E$
-In Algorithm 3, 2 and 4, we assume that the feature dimension for each tensor is all divisible by BLK.
-Notice that when accessing the re-index vector $v$, we may get value -1, since the workload for each expert is dynamic.
-In this case, we would skip this index, and remain zero for the temporary loading variable.
-We also provide the formulation for top-$k$ routing with our expert-specific operators, and take a comparison with top-$1$ routing.
-We denote the routing choice for top-$k$ as $\{\mathcal{R}_{i}(\bm{x})\}_{i=0}^{k-1}$, and other symbols keep consistent with Figure 5.
-We present the comparison on formulations for MoE forward and backward propagation in Table 5.
-Table 5: Comparison between top-$1$ and top-$k$ routing for the formulation of MoE forward and backward with our expert-specific operators.
-Stage Notation Layer Expert-Specific Formulation (top-$1$ routing) Expert-Specific Formulation (top-$k$ routing)
-Forward 1 1st MLP $\bm{y}_{1}=\textit{ESMM}(\bm{x},\bm{W}_{1},\bm{b}_{1},\mathcal{R}(\bm{x}))$$\{\bm{y}_{1}^{i}\}_{i=0}^{k-1}:\bm{y}_{1}^{i}=\textit{ESMM}(\bm{x},\bm{W}_{1},\bm{b}_{1},\mathcal{R}_{i}(\bm{x}))$
-2 Activation $\bm{y}_{2}=\mathcal{F}(\bm{y}_{1})$$\{\bm{y}_{2}^{i}\}_{i=0}^{k-1}:\bm{y}_{2}^{i}=\mathcal{F}(\bm{y}_{1}^{i})$
-3 2nd MLP $\bm{y}=\textit{ESMM}(\bm{y}_{2},\bm{W}_{2},\bm{b}_{2},\mathcal{R}(\bm{x}))$ $\bm{y}=\sum_{i=0}^{k-1}\textit{ESMM}(\bm{y}_{2}^{i},\bm{W}_{2},\bm{b}_{2},\mathcal{R}_{i}(\bm{x}))$
-Backward 4 2nd MLP $\frac{\partial\ell}{\partial\bm{b}_{2}}=\textit{ESS}(\frac{\partial\ell}{\partial\bm{y}},\mathcal{R}(\bm{x}))$ $\frac{\partial\ell}{\partial\bm{b}_{2}}=\sum_{i=0}^{k-1}\textit{ESS}(\frac{\partial\ell}{\partial\bm{y}},\mathcal{R}_{i}(\bm{x}))$
-5 $\frac{\partial\ell}{\partial\bm{W}_{1}}=\textit{ESTMM}(\bm{x},\frac{\partial\ell}{\partial\bm{y}_{1}},\mathcal{R}(\bm{x}))$ $\frac{\partial\ell}{\partial\bm{W}_{1}}=\sum_{i=0}^{k-1}\textit{ESTMM}(\bm{x},\frac{\partial\ell}{\partial\bm{y}_{1}^{i}},\mathcal{R}_{i}(\bm{x}))$
-6 $\frac{\partial\ell}{\partial\bm{y}_{2}}=\textit{ESMM}(\frac{\partial\ell}{\partial\bm{y}},\bm{W}_{2}^{T},\textit{null},\mathcal{R}(\bm{x}))$$\Big{\{}\frac{\partial\ell}{\partial\bm{y}_{2}^{i}}\Big{\}}_{i=0}^{k-1}:\frac{\partial\ell}{\partial\bm{y}_{2}^{i}}=\textit{ESMM}(\frac{\partial\ell}{\partial\bm{y}},\bm{W}_{2}^{T},\textit{null},\mathcal{R}_{i}(\bm{x}))$
-7 Activation $\frac{\partial\ell}{\partial\bm{y}_{1}}=\frac{\partial\ell}{\partial\bm{y}_{2}}\odot\mathcal{F}^{\prime}(\bm{y}_{1})$$\Big{\{}\frac{\partial\ell}{\partial\bm{y}_{1}^{i}}\Big{\}}_{i=0}^{k-1}:\frac{\partial\ell}{\partial\bm{y}_{1}^{i}}=\frac{\partial\ell}{\partial\bm{y}_{2}^{i}}\odot\mathcal{F}^{\prime}(\bm{y}_{1}^{i})$
-8 1st MLP $\frac{\partial\ell}{\bm{b}_{1}}=\textit{ESS}(\frac{\partial\ell}{\partial\bm{y}_{1}},\mathcal{R}(\bm{x}))$ $\frac{\partial\ell}{\bm{b}_{1}}=\sum_{i=0}^{k-1}\textit{ESS}(\frac{\partial\ell}{\partial\bm{y}_{1}^{i}},\mathcal{R}_{i}(\bm{x}))$
-9 $\frac{\partial\ell}{\partial\bm{W}_{1}}=\textit{ESTMM}(\bm{x},\frac{\partial\ell}{\partial\bm{y}_{1}},\mathcal{R}(\bm{x}))$ $\frac{\partial\ell}{\partial\bm{W}_{1}}=\sum_{i=0}^{k-1}\textit{ESTMM}(\bm{x},\frac{\partial\ell}{\partial\bm{y}_{1}^{i}},\mathcal{R}_{i}(\bm{x}))$
-10 $\frac{\partial\ell}{\partial\bm{x}}=\textit{ESMM}(\frac{\partial\ell}{\partial\bm{y}_{1}},\bm{W}_{1}^{T},\textit{null},\mathcal{R}(\bm{x}))$ $\frac{\partial\ell}{\partial\bm{x}}=\sum_{i=0}^{k-1}\textit{ESMM}(\frac{\partial\ell}{\partial\bm{y}_{1}^{i}},\bm{W}_{1}^{T},\textit{null},\mathcal{R}_{i}(\bm{x}))$
+Appendix B Algorithm Details
 Algorithm 1 Constructing re-index vector
@@
 Output: Tensor $v$ and $idx\_$
-Algorithm 2 Expert-specific summation
-[…]
-  Initialize: Output tokens $y$ with shape $(E,D)$ initialized with 0.
-  Parallel for $i=0$ to $E-1$ do
-   Parallel for $j$ in range $(0,D,\texttt{BLK})\ \textbf{do}$
-     $exp=\mathcal{R}[v[idx[i]]]$
-    Initialize zero tensor $c$ with shape $(1,\texttt{BLK})$
-[…]
-    for $k$ in range $(idx[i],idx[i+1],\texttt{BLK})\ \textbf{do}$
-     Parallel for $t=0$ to BLK  do
-      load $x_{sub}[t]=x[v[k+t],j:j+\texttt{BLK}]$
-     end for
-      $c=c+\sum\limits_{t=0}^{\texttt{BLK}-1}x_{sub}[t]$
-    end for
-   Write back: $y[exp,j:j+\texttt{BLK}]=c$
-   end for
-  end for
-  Output: Tensor $y$
 Algorithm 2 Expert-specific matrix multiplication
@@
   end for
+end for
+end for
+Output: Tensor $y$
+Algorithm 3 Expert-specific summation
+[…]
+Initialize: Output tokens $y$ with shape $(E,D)$ initialized with 0.
+Parallel for $i=0$ to $E-1$ do
+Parallel for $j$ in range $(0,D,\texttt{BLK})\ \textbf{do}$
+   $exp=\mathcal{R}[v[idx[i]]]$
+  Initialize zero tensor $c$ with shape $(1,\texttt{BLK})$
+[…]
+  for $k$ in range $(idx[i],idx[i+1],\texttt{BLK})\ \textbf{do}$
+   Parallel for $t=0$ to BLK  do
+    load $x_{sub}[t]=x[v[k+t],j:j+\texttt{BLK}]$
+   end for
+    $c=c+\sum\limits_{t=0}^{\texttt{BLK}-1}x_{sub}[t]$
+  end for
+Write back: $y[exp,j:j+\texttt{BLK}]=c$
 end for
@@
 Output: Tensor $y$
-Appendix B Experimental Details
-We provide the details of our CUDA program via enumerating the shape of the thread block and thread grid for expert-specific operators in a single MoE layer in Table 6.
+We present the algorithm details for re-index vector construction as well as the expert-specific operators, and taking top-1 routing as an example for illustration, shown in Algorithm 1, 2, 3 and 4.
+In Algorithm 1, we re-arrange the routing choice vector $\mathcal{R}$ into re-indexed token vector $v$, along with the token index starting vector $idx$, which satisfies $idx[0]=0$ and $idx[E]=N^{\prime}$.
+We provide the length and range for the vectors in Table 4.
+Table 4: Explanation for the auxiliary vectors.
+vector length range
+$\mathcal{R}$ $N$ $0\leq\mathcal{R}[i]<E$ for $0\leq i<N$
+$v$ $N^{\prime}$ $0\leq v[i]<N$ for $0\leq i<N^{\prime}$
+$idx$ $1+E$ $0\leq idx[i]<N^{\prime}$ for $0\leq i<1+E$
+In Algorithm 2, 3 and 4, we assume that the feature dimension for each tensor is all divisible by BLK.
+Notice that when accessing the re-index vector $v$, we may get value -1, since the workload for each expert is dynamic.
+In this case, we would skip this index, and remain zero for the temporary loading variable.
+We also provide the formulation for top-$k$ routing with our expert-specific operators, and take a comparison with top-$1$ routing.
+We denote the routing choice for top-$k$ as $\{\mathcal{R}_{i}(\boldsymbol{x})\}_{i=0}^{k-1}$, and other symbols keep consistent with Figure 2.
+We present the comparison on formulations for MoE forward and backward propagation in Table 5.
+Table 5: Comparison between top-$1$ and top-$k$ routing for the formulation of MoE forward and backward with our expert-specific operators.
+Stage Notation Layer Expert-Specific Formulation (top-$1$ routing) Expert-Specific Formulation (top-$k$ routing)
+Forward 1 1st MLP $\boldsymbol{y}_{1}=\textit{ESMM}(\boldsymbol{x},\boldsymbol{W}_{1},\boldsymbol{b}_{1},\mathcal{R}(\boldsymbol{x}))$$\{\boldsymbol{y}_{1}^{i}\}_{i=0}^{k-1}:\boldsymbol{y}_{1}^{i}=\textit{ESMM}(\boldsymbol{x},\boldsymbol{W}_{1},\boldsymbol{b}_{1},\mathcal{R}_{i}(\boldsymbol{x}))$
+2 Activation $\boldsymbol{y}_{2}=\mathcal{F}(\boldsymbol{y}_{1})$$\{\boldsymbol{y}_{2}^{i}\}_{i=0}^{k-1}:\boldsymbol{y}_{2}^{i}=\mathcal{F}(\boldsymbol{y}_{1}^{i})$
+3 2nd MLP $\boldsymbol{y}=\textit{ESMM}(\boldsymbol{y}_{2},\boldsymbol{W}_{2},\boldsymbol{b}_{2},\mathcal{R}(\boldsymbol{x}))$ $\boldsymbol{y}=\sum_{i=0}^{k-1}\textit{ESMM}(\boldsymbol{y}_{2}^{i},\boldsymbol{W}_{2},\boldsymbol{b}_{2},\mathcal{R}_{i}(\boldsymbol{x}))$
+Backward 4 2nd MLP $\frac{\partial\ell}{\partial\boldsymbol{b}_{2}}=\textit{ESS}(\frac{\partial\ell}{\partial\boldsymbol{y}},\mathcal{R}(\boldsymbol{x}))$ $\frac{\partial\ell}{\partial\boldsymbol{b}_{2}}=\sum_{i=0}^{k-1}\textit{ESS}(\frac{\partial\ell}{\partial\boldsymbol{y}},\mathcal{R}_{i}(\boldsymbol{x}))$
+5 $\frac{\partial\ell}{\partial\boldsymbol{W}_{1}}=\textit{ESTMM}(\boldsymbol{x},\frac{\partial\ell}{\partial\boldsymbol{y}_{1}},\mathcal{R}(\boldsymbol{x}))$ $\frac{\partial\ell}{\partial\boldsymbol{W}_{1}}=\sum_{i=0}^{k-1}\textit{ESTMM}(\boldsymbol{x},\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}},\mathcal{R}_{i}(\boldsymbol{x}))$
+6 $\frac{\partial\ell}{\partial\boldsymbol{y}_{2}}=\textit{ESMM}(\frac{\partial\ell}{\partial\boldsymbol{y}},\boldsymbol{W}_{2}^{T},\textit{null},\mathcal{R}(\boldsymbol{x}))$$\Big{\{}\frac{\partial\ell}{\partial\boldsymbol{y}_{2}^{i}}\Big{\}}_{i=0}^{k-1}:\frac{\partial\ell}{\partial\boldsymbol{y}_{2}^{i}}=\textit{ESMM}(\frac{\partial\ell}{\partial\boldsymbol{y}},\boldsymbol{W}_{2}^{T},\textit{null},\mathcal{R}_{i}(\boldsymbol{x}))$
+7 Activation $\frac{\partial\ell}{\partial\boldsymbol{y}_{1}}=\frac{\partial\ell}{\partial\boldsymbol{y}_{2}}\odot\mathcal{F}^{\prime}(\boldsymbol{y}_{1})$$\Big{\{}\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}}\Big{\}}_{i=0}^{k-1}:\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}}=\frac{\partial\ell}{\partial\boldsymbol{y}_{2}^{i}}\odot\mathcal{F}^{\prime}(\boldsymbol{y}_{1}^{i})$
+8 1st MLP $\frac{\partial\ell}{\boldsymbol{b}_{1}}=\textit{ESS}(\frac{\partial\ell}{\partial\boldsymbol{y}_{1}},\mathcal{R}(\boldsymbol{x}))$ $\frac{\partial\ell}{\boldsymbol{b}_{1}}=\sum_{i=0}^{k-1}\textit{ESS}(\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}},\mathcal{R}_{i}(\boldsymbol{x}))$
+9 $\frac{\partial\ell}{\partial\boldsymbol{W}_{1}}=\textit{ESTMM}(\boldsymbol{x},\frac{\partial\ell}{\partial\boldsymbol{y}_{1}},\mathcal{R}(\boldsymbol{x}))$ $\frac{\partial\ell}{\partial\boldsymbol{W}_{1}}=\sum_{i=0}^{k-1}\textit{ESTMM}(\boldsymbol{x},\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}},\mathcal{R}_{i}(\boldsymbol{x}))$
+10 $\frac{\partial\ell}{\partial\boldsymbol{x}}=\textit{ESMM}(\frac{\partial\ell}{\partial\boldsymbol{y}_{1}},\boldsymbol{W}_{1}^{T},\textit{null},\mathcal{R}(\boldsymbol{x}))$ $\frac{\partial\ell}{\partial\boldsymbol{x}}=\sum_{i=0}^{k-1}\textit{ESMM}(\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}},\boldsymbol{W}_{1}^{T},\textit{null},\mathcal{R}_{i}(\boldsymbol{x}))$
 Table 6: Shape of the thread block and thread grid for expert-specific operators in one MoE layer.
@@
 Operator Input Output Thread Block Thread Grid
-forward ESMM $\bm{x}$ $(N,D_{1})$ $\bm{y}$ $(N,D_{2})$ (WARP, TIMES) ($\lceil N^{\prime}/\texttt{BLK}\rceil$, $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
-$\bm{w}$ $(E,D_{1},D_{2})$
-$\bm{b}$ $(E,D_{2})$
-backward ESMM $\bm{x}$ $(N,D_{2})$ $\bm{y}$ $(N,D_{1})$ (WARP, TIMES) ($\lceil N^{\prime}/\texttt{BLK}\rceil$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
-$\bm{w}$ $(E,D_{2},D_{1})$
-ESS $\bm{x}$ $(N,D_{2})$ $\bm{y}$ $(E,D_{2})$ (WARP, TIMES) ($E$, $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
-ESTMM $\bm{x}_{1}$ $(N,D_{1})$ $\bm{w}$ $(E,D_{1},D_{2})$ (WARP, TIMES) ($\lceil D_{2}/\texttt{BLK}\rceil$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$, $E$)
-$\bm{x}_{2}$ $(N,D_{2})$
-ESFK $\bm{x}_{1}$ $(N,D_{1})$ $\bm{y}_{1}$ $(N,D_{1})$ (WARP, TIMES) ($E$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$, $\lceil N^{\prime}/\texttt{BLK}\rceil+\lceil D_{2}/\texttt{BLK}\rceil+$ $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
-$\bm{x}_{2}$ $(N,D_{2})$ $\bm{y}_{2}$ $(E,D_{2})$
-$\bm{w}_{1}$ $(E,D_{2},D_{1})$ $\bm{w}_{2}$ $(E,D_{1},D_{2})$
-We also provide the Python code for the proxy task we used to examine the computing capacity of the heterogeneous devices.
-We adopt a for loop composed of large matrix multiplications with the same scale as the test program, as provided below:
-import torch
-import time
-device = ’cuda’
-size = 2048
-times = 1024
-start_time = time.time()
-for j in range(times):
-mat1 = torch.randn(size, size, device=device)
-mat2 = torch.randn(size, size, device=device)
-y = torch.matmul(mat1, mat2)
-end_time = time.time()
-print(end_time - start_time)
-Appendix C Experimental Results
-We provide the exact values for both memory usage analysis and average latency analysis in Table 7 and 8, respectively.
+forward ESMM $\boldsymbol{x}$ $(N,D_{1})$ $\boldsymbol{y}$ $(N,D_{2})$ (WARP, TIMES) ($\lceil N^{\prime}/\texttt{BLK}\rceil$, $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
+$\boldsymbol{w}$ $(E,D_{1},D_{2})$
+$\boldsymbol{b}$ $(E,D_{2})$
+backward ESMM $\boldsymbol{x}$ $(N,D_{2})$ $\boldsymbol{y}$ $(N,D_{1})$ (WARP, TIMES) ($\lceil N^{\prime}/\texttt{BLK}\rceil$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
+$\boldsymbol{w}$ $(E,D_{2},D_{1})$
+ESS $\boldsymbol{x}$ $(N,D_{2})$ $\boldsymbol{y}$ $(E,D_{2})$ (WARP, TIMES) ($E$, $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
+ESTMM $\boldsymbol{x}_{1}$ $(N,D_{1})$ $\boldsymbol{w}$ $(E,D_{1},D_{2})$ (WARP, TIMES) ($\lceil D_{2}/\texttt{BLK}\rceil$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$, $E$)
+$\boldsymbol{x}_{2}$ $(N,D_{2})$
+ESFK $\boldsymbol{x}_{1}$ $(N,D_{1})$ $\boldsymbol{y}_{1}$ $(N,D_{1})$ (WARP, TIMES) ($E$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$, $\lceil N^{\prime}/\texttt{BLK}\rceil+\lceil D_{2}/\texttt{BLK}\rceil+$ $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
+$\boldsymbol{x}_{2}$ $(N,D_{2})$ $\boldsymbol{y}_{2}$ $(E,D_{2})$
+$\boldsymbol{w}_{1}$ $(E,D_{2},D_{1})$ $\boldsymbol{w}_{2}$ $(E,D_{1},D_{2})$
+Appendix C Experimental Details
+We provide the details of our CUDA program via enumerating the shape of the thread block and thread grid for expert-specific operators in a single MoE layer in Table 6.
+We also provide the PyTorch-style pseudocode for the proxy task we used to examine the computing capacity of the heterogeneous devices, as shown in Algorithm 5.
+We adopt a for loop composed of large matrix multiplications with the same scale as the test program.
+Algorithm 5 PyTorch-style pseudocode of MoE pipeline.
+1 import torch
+2 import time
+3
+4 device = ’cuda’
+5 size = 2048
+6 times = 1024
+7
+8 start_time = time.time()
+9 for j in range(times):
+10 mat1 = torch.randn(size, size, device=device)
+11 mat2 = torch.randn(size, size, device=device)
+12 y = torch.matmul(mat1, mat2)
+13 end_time = time.time()
+14
+15 print(end_time - start_time)
+Appendix D Experimental Results
+We provide the exact values for both memory footprint analysis and average latency analysis in Table 7 and 8, respectively.
 Specifically, for latency analysis, we provide the average value for each case with 0.5k, 1k, 1.5k and 2k total steps.
-We also visualize the validation accuracy of our method and Tutel under the same training time in Figure 25.
-We take Swin-MoE-Base with 4 experts and top-$1$ routing as the backbone network, training it on the homogeneous machine with the same batch size using Imagenet-1k dataset Deng et al.
-(2009).
-Since our HEXA-MoE can accelerate training process, its validation accuracy is much higher than baseline.
 Table 7: Memory analysis with Tutel, MegaBlocks and HEXA-MoE on Swin-Transformer-MoE benchmark (Base and Small).
 Experiments are conducted on 2 homogeneous GPUs with automatic mixed precision in PyTorch and batch size 40 for all the experiments.
-We set the number of global experts to 8, and record the average GPU memory consumption (GB) on each device.
+We set the number of global experts to 8, and record the average GPU memory footprint (GB) on each device.
 Method top-1 top-2 top-3 top-4 top-5 top-6 top-7 top-8
@@
 Ours (data-centric) 1.07 1.05 1.05 1.05 1.09 1.08 1.07 1.07
-Figure 25: Comparison on validation accuracy for Tutel and HEXA-MoE with the same training time. HEXA-MoE has a significant advantage over Tutel due to the highly improved efficiency on computation.